### v1

# Everything everywhere all at once, a probability-based enhanced sampling approach to rare events

Abstract
The problem of studying rare events is central to many areas of computer simulations. In a recent paper [Kang, P., et al., Nat. Comput. Sci. 4, 451–460, 2024], Kang, Trizio, and Parrinello (2024) we have shown that a powerful way of solving this problem passes through the computation of the committor function, and we have demonstrated how the committor can be iteratively computed in a variational way and the transition state ensemble efficiently sampled. Here, we greatly ameliorate this procedure by combining it with a metadynamics-like enhanced sampling approach in which a logarithmic function of the committor is used as a collective variable. This integrated procedure leads to an accurate and balanced sampling of the free energy surface in which transition states and metastable basins are studied with the same thoroughness. We also show that our approach can be used in cases in which competing reactive paths are possible and intermediate metastable are encountered. In addition, we demonstrate how physical insights can be obtained from the optimized committor model and the sampled data, thus providing a full characterization of the rare event under study. We ascribe the success of this approach to the use of a probability-based description of rare events.
I Introduction
Atomistic simulations hold the promise of allowing complex physicochemical processes to be studied with a minimum of theoretical assumptions. Frenkel and Smit (2001) However, one of the major limits of this approach is the gap between affordable simulation time and the time scale of many highly relevant phenomena, such as chemical reactions, protein folding, or crystallization. This is due to the presence of kinetic bottlenecks that slow down transitions between different states, which thus are called rare events. The relevance of this problem is witnessed by the ever-growing literature on the subject. Hénin et al. (2022)
In the last few decades, our group has contributed to this collective effort and very recently has proposed a new practical method to solve the rare event problem, which is based on computing the committor function $q(\mathbf{x})$ and using it as a sampling tool, which we summarize in Sec. III.1Kang, Trizio, and Parrinello (2024) We recall here that the committor is a function of the atomic coordinates $\mathbf{x}$ which, given two metastable states $A$ and $B$, is defined as the probability that a trajectory started at $\mathbf{x}$ ends in $B$ without having first passed by $A$Weinan and Vanden-Eijnden (2010) The committor function is arguably the most precisely defined way of describing rare events and is a quantity that remains well-defined even if the transition from $A$ to $B$ follows different competing pathways. As the committor has so many attractive features, much work has been done to determine it. Bolhuis, Dellago, and Chandler (2000); Ma and Dinner (2005); Berezhkovskii and Szabo (2005); Peters and Trout (2006); E. and Vanden-Eijnden (2006); Rotskoff, Mitchell, and Vanden-Eijnden (2022); Chen et al. (2023); Chen, Roux, and Chipot (2023); Li, Lin, and Ren (2019); Jung et al. (2023); Mitchell and Rotskoff (2024); He, Chipot, and Roux (2022); Khoo, Lu, and Ying (2019)
To evaluate $q(\mathbf{x})$ we have taken advantage of the variational principle of Kolmogorov, Kolmogoroff (1931) which states that if the boundary conditions $q({\mathbf{x}}_{A})=0$ and $q({\mathbf{x}}_{B})=1$ are satisfied for configurations $\mathbf{x}_{A}$ and $\mathbf{x}_{B}$ from the metastable basins, the committor can be computed by minimizing the functional:
$K[q(\mathbf{x})]=\langle|\nabla_{\mathbf{u}}q(\mathbf{x})|^{2}\rangle_{U(\mathbf{x})}$ (1)
where the average is over the Boltzmann distribution associated with the interatomic potential $U(\mathbf{x})$ at inverse temperature $\beta$, and $\nabla_{\mathbf{u}}$ denotes the gradient with respect to the mass-weighted coordinates.
In Ref. 1, we computed the variational minimum of $K[q(\mathbf{x})]$ using a self-consistent iterative procedure that to be started, only requires that the initial and final states are known. Key to the success of this approach has been the use of a bias potential that is a functional of the committor
$V_{K}(\mathbf{x})=-\frac{1}{\beta}\log(|\nabla q(\mathbf{x})|^{2})$ (2)
which is able to focus sampling on the otherwise elusive transition state region (see Fig. 1). This approach has proven to be highly successful in finding and extensively sampling the transition state ensemble (TSE). However, additional simulations were needed to evaluate the system free energy since $V_{K}(\mathbf{x})$ disproportionately favors sampling of the transition state ensemble region. This additional calculation can be performed, for instance, using the data collected during the iterative process to generate an efficient collective variable Bonati, Piccini, and Parrinello (2021); Ray, Trizio, and Parrinello (2023) (CV) and then performing a metadynamics-like calculation using this CV. Laio and Parrinello (2002); Invernizzi and Parrinello (2020)
Figure 1: Schematic representation of enhanced sampling framework. A. Relationship between the committor function $q$ and the $z$ CV. In the model, a set of input descriptors $\mathbf{d}(\mathbf{x})$ is combined using a neural network into a one-dimensional latent space $z$. The final output that, upon optimization, represents the committor function is obtained by applying a sigmoid-like activation function $\sigma$ on $z$. As a consequence, $z$ and $q$ encode the same information, but the $z$ space is more spread and smooth with respect to the sharper $q$ space, and it is thus more suitable to be used in an enhanced sampling context. B. Different scenarios from the energy and probability points of view (top and bottom rows, respectively) for a double-well model potential with two states, A and B. Left to right, we show the unbiased scenario, a biased scenario where the metadynamics-like OPES bias fills the basins to promote transition, a biased scenario where the Kolmogorov bias ($V_{K}$) stabilizes the transition state (TS), a biased scenario in which the combined bias OPES+$V_{K}$ allows extensively sampling the metastable and transition states and promoting transitions between them.
However, in principle, such a post-processing search for a good CV should not be necessary since the committor has been argued to be the best possible one-dimensional reaction coordinate. Berezhkovskii and Szabo (2005); Peters and Trout (2006); Ma and Dinner (2005); Li and Ma (2014); He, Chipot, and Roux (2022) Unfortunately, directly using $q(\mathbf{x})$ as a CV is unpractical, if not impossible, in an enhanced sampling context. Indeed, in the basins, where $q(\mathbf{x})$ is either $\approx 0$ ($\mathbf{x}\in A$) or $\approx 1$ ($\mathbf{x}\in B$), different configurations are distinguished only by very tiny numerical variations of $q(\mathbf{x})$, thus leading to numerical difficulties. Furthermore, another source of numerical issues comes from the very sharp behavior of $q(\mathbf{x})$ in the transition state region, Khoo, Lu, and Ying (2019) further complicating its use as a CV (see Fig. 1A and SI, Fig. S1).
However, a convenient solution can be found if we recall that in Ref. 1 $q(\mathbf{x})$ was expressed as a neural network whose output $z(\mathbf{x})$ was fed to a step-like activation function $\sigma$ to facilitate imposing the right functional form (see Fig. 1A)
$q(\mathbf{x})=\sigma\big{(}z(\mathbf{x})\big{)}$ (3)
The output of the network $z(\mathbf{x})$ encodes the same information as $q(\mathbf{x})$ but is smoothly varying and can be used in practice as an efficient CV. This observation is crucial to the success of the improved approach we present here, which makes the method of Ref. 1 an effective and semi-automatic way to perform free energy calculations. Indeed, we do not wait for the end of the iterative process to perform an additional free energy calculation, but we add to $V_{K}$ an On-the-fly Probability Enhanced Sampling Invernizzi and Parrinello (2020) (OPES) bias that uses $z(\mathbf{x})$ as a CV in the iterative procedure itself and it’s aimed at filling the energy landscape to promote transitions (see Fig. 1B). We shall refer to this combined bias as OPES+$V_{K}$, and we show that this, alongside a few technical improvements, accelerates the convergence of the iterative procedure and leads to a balanced sampling, in which configurations belonging to the metastable states and to the transition state ensemble are equally well sampled, and no additional calculations are needed to obtain a converged free energy estimate (see Fig. 1B).
For illustrative purposes, we first demonstrate the workings of our approach on some simple but informative systems like some two-dimensional models and the classical example of alanine dipeptide in a vacuum. We then move to more complex systems, studying the folding of the chignolin protein in water and the interaction between an organic ligand and a calyxarene, which is a simplified but non-trivial model for drug-protein interaction. In this last example, different binding pathways are possible, and the ligand visits several intermediate states on its way to the binding pocket. This demonstrates the generality of our approach, which does not require assuming that there is only one reaction pathway. A full description of the methods, a summary of the procedure, and other technical details are provided in Sec. III and in the SI.
II Results
II.1 Müller potential
Figure 2: Müller-Brown potential. A. Scatter plot of the sampling under the coupled action of the OPES bias along the $z$ CV and the TS-oriented Kolmogorov bias $V_{K}$ at convergence after two iterations of our protocol. The isolines outline the energy levels of the potential, and points sampled starting the simulation from basins A and B are reported in blue and red, respectively. To demonstrate convergence, we report the physically relevant part of the $q\sim 0.5$ isoline for a numerical reference for the committor (grey line), and the one learned from our model (yellow). The outputs of the same two functions on all the sampled points are compared in the inset, in which the yellow dashed line gives the equivalence line. B. Distribution of the sampled configurations along the $z$ CV for simulations driven by OPES+$V_{K}$ (blue curve) and only by OPES (red curve). The black dashed lines indicate the relative peaks of the two distributions. C. Free energy estimate along $z$ obtained from the two simulations reported in B and from a numerical reference (grey dotted line). The black dashed lines are consistent with those reported in B.
We start discussing the merits of our new approach on the simple but instructive example of a single particle diffusion on the Müller-Brown potential. As in our previous work, Kang, Trizio, and Parrinello (2024) we use the $x$ and $y$ coordinates of the particle as descriptors, thus allowing a fair comparison between the two methods.
Performing OPES+$V_{K}$ simulations allows covering the whole relevant phase space from the metastable states through the TS, as shown in Fig. 2A. Moreover, to achieve this better sampling, it is not necessary to wait until a converged model is available, as good sampling is obtained even from the very first iteration (see SI, Fig. S2) in which the model is not much more than a classifier trained to satisfy the boundary conditions $q({\mathbf{x}}_{A})=0$ and $q({\mathbf{x}}_{B})=1$ (see Sec. III.2). However, this result should come as little surprise, as machine-learning classifier-like CVs have already proved effective in driving enhanced sampling simulations with OPES. Bonati, Rizzi, and Parrinello (2020); Trizio and Parrinello (2021); Ray, Trizio, and Parrinello (2023) In contrast, using only $V_{K}$ in Ref. 1, the sampling was unbalanced toward the TS region and often heavily dependent on the basin from which the simulation was started. Another advantage in terms of sampling comes from the action of the Kolmogorov bias $V_{K}$, which favors extensive sampling of the TS region. This is evident if one compares the distribution along $z$ of the points sampled with a standard OPES simulation and one performed by combining OPES and $V_{K}$ (see Fig. 2B). From the two distributions, it can be clearly appreciated the different emphasis that the two simulations put on the TS region. When only OPES is used, most of the time is spent in the metastable basins. On the other hand, with the combined bias, a large number of TSE configurations is harvested. Such configurations are precious not only for better investigating the TS properties but also for effectively applying the variational principle to optimize the committor.
Figure 3: Alanine dipeptide. Top row: normalized densities in the plane of $\varphi$ and $\theta$ torsional angles of the sampled configurations (A) under the joint action of the OPES bias on the $z$ CV and the TS-oriented Kolmogorov bias $V_{K}$ and of the corresponding reweighted Kolmogorov distribution $p_{K}$ (B). The white isolines provide the reference free energy levels, whereas the magenta dotted line in C highlights the relation between $\varphi$ and $\theta$ for TSE configurations proposed in Ref. 5. Bottom row: comparison of the distribution of sampled configurations (C) and free energy estimates (D) along $\varphi$ with different enhanced sampling setups: OPES on $\varphi$ and $\psi$ (grey), OPES on $z$ (red), and OPES+$V_{K}$ on $z$ (blue, same data as A).
In addition, a second advantage of this approach is related, besides the quantity of the sampling, to its quality. Thanks to the static nature of $V_{K}$ and the robust OPES framework, Invernizzi and Parrinello (2022) which is designed to quickly converge to a quasi-static regime, very reliable weights can be assigned to the sampled points, thus allowing for a proper reweighting to the true Boltzmann distribution, which, again, has several implications. On the one side, this allows the recovery of the free energy surface (FES), as we show in Fig. 2C. There, we report the free energy profiles along $z$ obtained by reweighting the same two trajectories considered above, compared with an analytical reference. The three curves are almost indistinguishable, thus proving the quality of the reweighting procedure and the (non-obvious) absence of detrimental effects of the $V_{K}$ bias on such free energy calculations.
On the other side, having such reliable weights also for the points in the metastable regions facilitates the convergence of the iterative procedure, at variance with Ref. 1. Indeed, even for this simple example, this approach allows converging the estimate of the committor to the numerical reference almost perfectly in just a couple of iterations (see inset of Fig. 2A).
II.2 Alanine Dipeptide
In most publications devoted to rare event sampling, it is traditional to illustrate the capability of a new method on the conformational equilibrium of alanine dipeptide, and here, we shall stick to this tradition. Moreover, having studied alanine also in our previous work, Kang, Trizio, and Parrinello (2024) we are in the position to asses the improvement in convergence speed due to our improved protocol. There, a good guess for the committor function was obtained from the $6^{\text{th}}$ iteration of the procedure, whereas here, it was already available with half of the iterations (see SI, Table S2). This allowed us not only to efficiently sample the whole relevant phase space (see Fig. 3A) but also to properly identify the TSE (Fig. 3B), which is known to exhibit a linear relation between $\varphi$ and $\theta$.  Bolhuis, Dellago, and Chandler (2000) It is also worth noting that one can still estimate the Kolmogorov distribution $p_{K}$ by reweighting the OPES+$V_{K}(\mathbf{x})$ sampled configurations only with the OPES part of the bias. As discussed in Ref.1, this allows defining in a precise way the TSE.
Also in this case, it is instructive to compare the distribution of points sampled using different bias potentials. In Fig. 3C, we compare the outcome of three different OPES simulations, in one, we use the traditional $\varphi$ and $\psi$ as CVs, in the second the $z(\mathbf{x})$ is used as CV, and finally in the third the action of the $z(\mathbf{x})$ OPES bias is boosted by the Kolmogorov bias $V_{K}(\mathbf{x})$. As seen for the Müller potential, combining the $V_{K}$ potential with OPES systematically enhances the sampling of the TSE while still guaranteeing a thorough sampling of the metastable states. From such data, we can thus recover a correct estimate of the free energy of the system, which is indistinguishable not only from that obtained in the normal OPES run on $z$ but also from the reference simulation using $\varphi$ and $\psi$ (see Fig. 3D), thus increasing our confidence in the method.
II.3 Double Path Potential
Figure 4: Asymmetric double path potential. A. Contour plot of the learned committor function superimposed to the isolines of the potential. B. Kolmogorov distribution $p_{K}$ computed from a simple reweighting of the sampled trajectory by the OPES bias only. C. Two-dimensional free energy surface recovered by reweighting the sampled trajectories by the total effective bias OPES+$V_{K}$.
Another situation that can occur in practice and which is not straightforward to deal with using other methodologies is the one in which the reaction can follow different pathways Vlugt and Smit (2001); Borrero and Dellago (2016); Bolhuis and Csányi (2018); Mandelli, Hirshberg, and Parrinello (2020); Capelli, Carloni, and Parrinello (2019). A simplified and yet instructive model is the one depicted in Fig. 4, where the system can go from state $A$ to state $B$ following two different pathways, with the upper path characterized by a larger free energy barrier.
With our approach, both paths are sampled starting from the first iteration, thus allowing for a quick convergence of the procedure (see SI, Fig. S4). It is then instructive to focus on the final results reported in Fig. 4. In panel A, we report the converged committor function for the system, showing not only that both paths can be handled at the same time but also that the committor varies more abruptly in the case of the higher barrier, whereas it is more spread for the favored channel. This behavior provides a non-trivial example of why our choice to use the Kolmogorov distribution Kang, Trizio, and Parrinello (2024) $p_{K}$ (see Eq. 7) to identify TSE is to be preferred over the conventional $q\simeq 0.5$ definition. If one considers the $p_{K}$ for this system (see Fig. 4), one can readily see that most of the contribution to the transition flux comes from the lower channel, and only a minor portion goes through the other. In contrast, the $q\simeq 0.5$ criterion would have given the same importance to the two paths and also to much higher energy regions that were actually never sampled, not even in biased simulations (see SI, Fig. S4). The predominance of the lower path is also reflected by the distribution of the sampled points, which are much denser there (see SI, Fig. S4). Nevertheless, after reweighting, an almost perfect 2D FES for the system, including precise estimates of the barriers, can be recovered effortlessly, as shown in Fig. 4C.
It is worth noticing that all these results were obtained without any prior information on the system if not knowledge of the initial and final metastable states. From this little amount of information, through the iterative procedure and thanks to the combined action of the OPES+$V_{K}$, we could not only blindly discover the two possible paths but also fully characterize them. Another more realistic example of this situation will also be shown in the following.
[…]
II.4 Chignolin: Protein Folding
A more challenging test for our method is the study of the folding of the solvated chignolin protein into a stable hairpin structure. Conveniently, for this system, results from long unbiased MD simulations are also available, Lindorff-Larsen et al. (2011) providing a reference for the folding energy and the overall dynamics of the process. In our previous work, Kang, Trizio, and Parrinello (2024) we already studied this system, characterizing the TSE committor and identifying two possible transition routes. There, we used as descriptors the 45 distances between the alpha carbon atoms, which were able to describe well the protein backbone. Here, we make use of the greater efficiency of our approach to enlarge the number of descriptors, using the 210 distances selected in Ref. 19 that also include the interaction of the backbone with the side chains. As discussed in Ref. 1, better variational flexibility typically leads to more accurate results and to a lower value of the variational bound $K_{m}$. Here, in fact, on the same dataset, the minimum goes from the $K_{m}\approx 19$ to the much lower bound of $K_{m}\approx 2.2$ (see also SI, Sec S4).
[…]
In Ref. 1, we have shown that the properties of the transition state ensemble could be analyzed in great and illuminating detail with a combination of k-medoid clustering Schubert and Lenssen (2022) and feature relevance analysis. Pizarroso, Portela, and Muñoz (2022) Here, as we get a thorough sampling of the entire reaction path, we extend this analysis to understand the full folding mechanism.
To do so, we take advantage of the fact that $q(\mathbf{x})$, or equivalently, for simplicity, the smoother $z(\mathbf{x})$, provide a natural ordering in the configurations sampled, that can be used to order them as one goes from $q(\mathbf{x}\in A)=0$ to $q(\mathbf{x}\in B)=1$. Thus, we sliced the whole trajectory into 10 bins along the whole $z$ range (i.e., $-6.5<z<6.5$) and into 10 more focused on the TS region (i.e., $-1.5<z<1.5$), and for each, we identified the most representative configurations through the k-medoid analysis. Schubert and Lenssen (2022) Then, we performed a LASSO-based Novelli et al. (2022) feature relevance analysis to identify which descriptors were more relevant for each bin, revealing that the orientation of the side chains does play a role in the formation of the hydrogen bond network that holds the protein together and regulate steric effects. In Fig. 5, we show the superimposition of several configurations belonging to the most representative bins across the (un)folding process.
A first general observation is that, as the folding proceeds, the entropy within the bins decreases, and the corresponding configurations are more similar to each other. Interestingly, an explanation can be found guided by the LASSO results (see SI, Table S5), which focus the attention on the sidechains of residues Tyr2 and Trp9 in the folded configurations (see Fig. 5 1 and 2). Indeed, moving from frame 1 to 2, the stable folded structure starts to be dismantled due to the misalignment of the side groups of Tyr2 and Trp9, which destroys the hydrophobic $\pi$-$\pi$ interaction between their aromatic rings and owing to the rotation of the side chain of Tyr1, which induces the breaking of the H-bond that locks the end of the main chain.
[…]
Finally, once at least part of the H-bond network is released, the whole structure of the protein is not stable anymore and thus progressively unfolds, completely losing the characteristic hairpin structure (6 and 7).
[…]
II.5 Calixarene: Host-Guest systems
The final test for our method is the study of the interaction of the G2 ligand with an octa-acid calixarene host (OAMe) from the SAMPL5 challenge, Yin et al. (2017) which, although simple, displays some of the key features of the biologically relevant protein-ligand systems.
Previous work Rizzi et al. (2021) has studied in detail this system by applying OPES on two independent CVs and using a funnel restraint to limit the unbound state diffusion. Limongelli, Bonomi, and Parrinello (2013) One CV was the projection of the ligand center of mass on the binding axis ($h$), and the other was a machine-learning CV based on a classification criterion Bonati, Rizzi, and Parrinello (2020) that accounted for the role of water in the process, as it took as inputs the water coordination numbers of a curated set of points along $h$ and on the ligand molecule. Here, as inputs for a single committor model, we use both types of information to obtain a simple and complete description of the whole process. More in detail, we used the same water coordination numbers and, rather than the $h$ projection, a set of distances between the ligand molecule and the binding pocket to have more flexibility in also accounting for the relative orientation of the ligand.
[…]
Moreover, our committor-based simulations allowed us to characterize different reaction channels in a single calculation, which we depict in Fig. 6 and can be identified by visualizing the projection of the Kolmogorov ensemble in the plane defined by $h$ and the water coordination number inside the cavity $V_{2}$ (see SI Fig. S9). As one could expect, the TS configurations are mostly determined by the distance of the ligand from the host molecule, as proved also by the feature relevance analysis (see SI, Fig. S6), which is sensibly shortened with respect to the fully unbound state. However, from $p_{K}$, it can also be seen that there are two contributions to the TSE, differing in the behavior of water and, similar to what is discussed in Sec. II.3, one dominates over the other.
[…]
Figure 7: Convergence of $\boldsymbol{\Delta}\mathbf{G}$ Convergence with simulation time of the estimate for the folding energy of chignolin (A) and the binding energy of the G2 ligand to the OAMe octa-acid guest (B). The average estimates from 3 independent simulations for each system are reported as a blue solid line, whereas the uncertainty, computed as the standard deviation over the three replicas, is depicted as a shaded blue region. The reference values are provided as grey dashed lines, and the $0.5\,k_{B}T$ interval around the reference is marked by grey dotted lines. For chignolin, that is the unbiased estimate from Ref. 32, for the calixarene, an estimate obtained using the enhanced sampling setup of Ref. 37.
III Methods
In the following sections, we first provide a general background on the committor and the variational principle it obeys, and we summarize the concepts of the Kolmogorov bias and the related ensemble we have introduced in Ref.1 (Sec. III.1). We then discuss the improved iterative protocol for the self-consistent determination of the committor function (Sec. III.2) and its use to converge free energy calculations (Sec. III.3), highlighting the advancements from our previous work. Kang, Trizio, and Parrinello (2024) In the last section, we schematically summarize the whole iterative procedure to help contextualize all the details (Sec. III.4).
III.1 Background
The committor function The committor function $q(\mathbf{x})$ was introduced by Kolmogorov for the study of rare events involving transitions between two metastable states $A$ and $B$Kolmogoroff (1931) Such a function is defined as the probability that a trajectory started at configuration $\mathbf{x}$ ends up in $B$ before first passing by $A$ Weinan and Vanden-Eijnden (2010) and is thus believed to be the best possible formal description for a rare event process. Berezhkovskii and Szabo (2005); Li and Ma (2014); Ma and Dinner (2005); He, Chipot, and Roux (2022) Unfortunately, determining the committor function is a far from trivial challenge.
In the transition path sampling literature, Dellago, Bolhuis, and Geissler (2006); Jung et al. (2023); Bolhuis et al. (2002) the committor value for a specific configuration $\mathbf{\tilde{x}}$ is evaluated performing the so-called committor analysis, in which many independent trajectories are started from $\mathbf{\tilde{x}}$ to statistically evaluate $q(\mathbf{\tilde{x}})$ based on the outcome of such simulations. Despite its appealing simplicity, this approach presents some serious drawbacks that limit its viability. For example, its computational cost can easily become very heavy as the number of selected $\mathbf{\tilde{x}}$ configurations grows, also because many simulations have to be performed for each $\mathbf{\tilde{x}}$ to obtain meaningful statistics. More critically, it is also known that when one tackles more complex systems, the results of such an analysis can be strongly affected by the criteria used to determine whether a trajectory is committed to either basin A or B. Lazzeri et al. (2023)
However, Kolmogorov also showed a theoretical approach to compute the committor function for a generic system governed by an interatomic potential $U(\mathbf{x})$ at inverse temperature $\beta$. This is rigorously derived under the hypothesis of overdamped Langevin dynamics, but with a few reasonable assumptions, it can also be extended to the general case. Weinan and Vanden-Eijnden (2010) Conveniently, such an approach can be equivalently formulated as a differential or a variational problem.
The first one is based on a set of partial differential equations that obey the boundary conditions $q(\mathbf{x}_{A})=0$ and $q(\mathbf{x}_{B})=1$ on the starting and final points of the transition $\mathbf{x}_{A}$ and $\mathbf{x}_{B}$. Unfortunately, a practical application of this approach is unfeasible, as its equations are impossible to solve if not for a few trivial low-dimensional toy models. The second approach, on the other hand, finds an equivalent solution following a variational principle that obeys the same boundary conditions. In practice, this amounts to minimizing the functional of the committor
[…]
As it often happens with variational problems, applying this second approach is less demanding, making it a viable option. The functional in Eq. 4 can indeed be written equivalently as the expectation value of the square modulus of the committor gradients $|\nabla_{\mathbf{u}}q(\mathbf{x})|^{2}$, which is related to the reaction rate $\nu_{R}$, and the average is over the Boltzmann ensemble
$K[q(\mathbf{x})]=\langle|\nabla_{\mathbf{u}}q(\mathbf{x})|^{2}\rangle_{U(\mathbf{x})}$ (5)
and thus, it can be estimated through sampling.
However, even this approach still presents some serious challenges. In a rare event scenario, the committor is indeed almost constant in the metastable basins, i.e., $q\simeq 0$ in A and $q\simeq 1$ in B, and smoothly interpolates between these two values in the transition region. As a consequence, the gradients that appear in Eq. 5 are strongly localized on the transition region, which, unfortunately, is exactly the region that is difficult to sample in simulations. To bypass this sampling difficulty, in Ref. 1, we introduced the Kolmogorov bias potential, a new enhanced sampling protocol that, by design, focuses the sampling on the TS, which we introduce in the following section.
Sampling the Kolmogorov ensemble In a rare event scenario, metastable states are separated by large free energy barriers that hinder the transitions between them. The set of high-energy configurations that are found at the top of the saddle points associated with such barriers is usually referred to as the transition state ensemble (TSE). Knowledge of the TSE is highly desired as it can provide precious insights into reaction mechanisms and rates, but unfortunately, sampling the TSE with MD simulations is typically an issue.
Such configurations are indeed completely out of reach in standard MD and are difficult to sample even when conventional metadynamics-like enhanced sampling methods are applied. Laio and Parrinello (2002); Invernizzi and Parrinello (2020) Indeed, in that case, the bias potential added to the system aims at filling the metastable basins to promote transitions between them by lowering the relative free energy barriers. However, in the biased energy landscape, the TS region still remains a local maximum whose sampling is thus unfavored compared to the metastable basins (see Fig. 1). In addition, the quality of the TSE sampling with such calculations also depends on the quality of the used CV.
Ideally, if one wants to study the TS, one would like to overturn this scenario, thus building a bias potential that, by design, focuses the sampling on the TS rather than wasting precious simulation time in the metastable basins. However, this sounds like a chicken-and-egg problem, as to build such a bias, we need to locate the TS, which is the very reason why we want to build the bias. Remarkably, a solution to this dilemma can be obtained by taking inspiration from the committor itself.
In Ref. Kang, Trizio, and Parrinello (2024), we introduced the Kolmogorov TS-oriented bias potential
$V_{K}(\mathbf{x})=-\frac{\lambda}{\beta}\log(|\nabla q(\mathbf{x})|^{2})$ (6)
where $\lambda\simeq 1$ is a parameter used to scale the magnitude of the bias. $V_{K}$ depends on the gradients of the committor function, which, as discussed in the previous section, are localized on the TS by construction. As a consequence, in the resulting biased energy landscape $U_{K}(\mathbf{x})=U(\mathbf{x})+V_{K}(\mathbf{x})$, the TS region is stabilized over the rest of the phase space, favoring its sampling (see Fig. 1).
Starting from $U_{K}$, we can define the Kolmogorov probability distribution
[…]
in which only those configurations that actually contribute to the reaction rate have a non-negligible contribution. Indeed, in such a distribution, the relative weight of the configurations is determined not only by the probability that state $\mathbf{x}$ is sampled, which is proportional to the Boltzmann weight $e^{-\beta U(\mathbf{x})}$, but also by the contribution that successful reactive trajectories passing by $\mathbf{x}$ bring to the reaction rate, which is expressed by the $e^{-\beta V_{K}(\mathbf{x})}=e^{-\log(|\nabla q(\mathbf{x})|^{2})}=|\nabla q(\mathbf{x})|^{2}$ term. For this reason, we proposed using the $p_{K}(\mathbf{x})$ to formalize an improved definition of the TSE, thus extending the conventional definition that includes all those configurations found on the $q\simeq 0.5$ isosurface, regardless of their actual probability of being sampled in the true Boltzmann distribution.
Self-consistent approach for the committor In Ref. 1, we have shown how the variational principle and the sampling of the Kolmogorov ensemble introduced above can be combined in a self-consistent iterative procedure to obtain the committor function with the help of machine learning. In the iterative procedure, we alternate training and sampling cycles until convergence is reached. The first are aimed at improving a neural-network-based variational parametrization of the committor, and the second generate new and progressively better data to learn from under the effect of the TS-oriented Kolmogorov bias. In the following sections, we present an improved version of the procedure that not only allows for a faster convergence through the iterations but also the calculation of the free energy profile of the process.
III.2 Machine learning the committor function
In the method presented in Ref. 1, the committor function is represented as a neural network (NN) $q_{\theta}(\mathbf{x})=q_{\theta}(\mathbf{d}(\mathbf{x}))$ with learnable parameters $\theta$ that takes as inputs a set of physical descriptors $\mathbf{d}(\mathbf{x})$ (see Fig. 1). Such descriptors are functions of the atomic coordinates and can be, for instance, distances or coordination numbers. The optimization of $q_{\theta}^{n}(\mathbf{x})$, at iteration $n$, is based on the minimization of an objective function $L$ that formalizes the variational principle of Eq. 5 and its boundary conditions. In fact, the loss function presents two terms $L=L_{v}+\alpha L_{b}$, relatively scaled by the hyperparameter $\alpha$. The $L_{v}$ term derives from the variational principle and reads
$L_{v}=\frac{1}{N^{n}}\sum_{i}^{N^{n}}w_{i}|\nabla_{\mathbf{u}}q(\mathbf{x}_{i})|^{2}$ (8)
where the weights $w_{i}$ reweigh each configuration $\mathbf{x}_{i}$ to it natural Boltzmann weight. The $L_{b}$ term enforces the boundary conditions $\mathbf{x}_{A}=0$ and $\mathbf{x}_{B}=1$ as
$L_{b}=\frac{1}{N_{A}}\sum_{i\in A}^{N_{A}}(q(\mathbf{x}_{i}))^{2}+\frac{1}{N_{B}}\sum_{i\in B}^{N_{B}}(q(\mathbf{x}_{i})-1)^{2}$ (9)
In Ref. 1, the $L_{b}$ term was evaluated on a labeled dataset of $N_{A}$ and $N_{B}$ configurations collected at the first iteration by running short, unbiased MD simulations in the two basins. On the other hand, the $L_{v}$ term was computed on the whole dataset available, thus including both the initial unbiased and biased $N^{n}$ configurations from later iterations. This was helpful, as the sampling cycles under the action of the $V_{K}$ bias were focused on the TS region but not optimal, as the unbiased configurations were incorrectly weighted in the dataset. This issue made it necessary to perform more iterations or correct the corresponding weights (i.e., $w_{i}$ when $i\in A,B$) by an estimate of the free energy difference between the two basins, which is not always readily available.
Our new approach gets rid of this slowdown, as at each sampling iteration, we can effectively visit not only the TS region but also both metastable states (see Sec. III.3). In this way, the new configurations cover the whole phase space, and after proper reweighting, the weights $w_{i}$ automatically follow the real Boltzmann distribution. As a consequence, we can thus avoid evaluating the $L_{v}$ term on the unbiased labeled configurations, making the whole procedure more precise and faster. It should also be noted that, in the first iteration, when only unbiased data are available, the contribution from the $L_{v}$ term was almost null also in the original formulation due to the localization of the gradients $|\nabla q(\mathbf{x})|^{2}$ discussed in Sec. III.1. A second consequence of the higher quality of the data thus generated is a more data-efficient protocol. Indeed, as soon as the enhanced sampling becomes effective (i.e., multiple transitions can be observed), to optimize the parameters $\theta$ at each iteration, it is sufficient to use configurations and weights generated in the previous iteration or, at most, from the previous two.
III.3 Extensive sampling along a committor-based CV
As already anticipated in the introduction, we bring two conceptual improvements to the original recipe of Ref. 1. First, we define a physics-informed collective variable (CV) $z(\mathbf{x})$ that is based on the committor and is effective when used with metadynamics-like enhanced sampling approaches. Second, we improve such approaches by combining them with the Kolmogorov bias potential to favor the TS sampling.
Even if the committor has often been thought to be an ideal CV, Berezhkovskii and Szabo (2005); Li and Ma (2014); Ma and Dinner (2005); He, Chipot, and Roux (2022) we already pointed out in the introduction that its straightforward application in an enhanced sampling context is far from optimal due to the high degeneracy of the metastable states along such a variable, as $q(\mathbf{x_{A}})\simeq 0$ and $q(\mathbf{x_{B}})\simeq 1$, and to its sharp behavior in the transition region. This would suggest that a smoother modification of the committor might be an effective CV. A practical way of simply and economically achieving this result comes at zero cost from the NN-based parametrization of the committor proposed in Ref. 1. There, as we have seen in Sec. III.2 and depicted in Fig. 1, the committor is represented as a NN, and a sharp sigmoid activation function $\sigma$ is applied to the output layer of such a network to enforce the correct functional form to the learned function, i.e., $q(\mathbf{x})=\sigma(z(\mathbf{x}))$. It is thus wiser to employ as a CV, not the activated final committor function, but rather the output $z(\mathbf{x})$ of the NN. In fact, this still preserves the physical information encoded in the committor function, but it is also suitable to be used in an enhanced sampling context as it still has finite gradients in the metastable regions, and it is not as sharp in the TS region.
As one can expect, similar numerical issues could also affect the bias potential $V_{K}$ that, in Ref. 1, was computed as a function of the gradients of the committor function $\nabla q(\mathbf{x})$. In fact, this choice is not optimal in practice as, due to the sharp behavior of $q(\mathbf{x})$, such a computation can easily lead to extremely small values below machine precision in the basins where $q(\mathbf{x})$ is flat (see SI, Fig. S1), thus making the bias almost ineffective there. Conveniently, by taking advantage of the change of variable $q\rightarrow z$ and with the help of some elementary algebra (see SI, Sec. S1), one can bypass this issue, writing the bias potential as a function of the smoother and more numerically stable gradients of $z$.
Like any CV, $z$ can be used, in principle, with any CV-based method. However, here, we opt for the On-the-fly Probability Enhanced Sampling (OPES) method Invernizzi and Parrinello (2020), a recent evolution of metadynamics Laio and Parrinello (2002) that arguably represents the state of the art in this field. We already pointed out in Ref. 1 that, in metadynamics-like frameworks, the metastable basins are filled with an external bias potential to reduce the height of free energy barriers and promote transitions, but the TS-region is still found at a local maximum, whose sampling is thus unfavored (see Fig. 1).
Here, we propose to remove this limitation by adding a Kolmogorov bias $V_{K}$ to such a biasing scheme. In fact, this will stabilize the TS, turning it into a minimum that will eventually be filled by the original OPES bias $V_{\text{OPES}}$ alongside the true metastable states. In practice, this amounts to driving the system with an effective bias potential $V_{\text{eff}}=V_{K}+V_{\text{OPES}}$. In this way, one can promote frequent reactive events while increasing the time spent in the usually elusive transition region. As a result, one can thus fully characterize the rare event by harvesting in the same simulation a wealth of precious TSE configurations and obtaining an accurate estimate of the free energy surface (FES). This can be recovered with a simple reweighting by the total effective bias potential according to
$w^{n}_{i}=\frac{e^{\beta V_{\text{eff}}^{n}(\mathbf{x}_{i})}}{\langle e^{\beta V_{\text{eff}}^{n}(\mathbf{x})}\rangle_{U^{n}_{\text{eff}}}}$ (10)
which benefits from the fast-converging OPES framework and the static nature of the Kolmogorov bias.
As one can expect, the improved sampling of this new approach is also beneficial to the iterative procedure as it greatly simplifies the dataset construction, exploiting the best from the two biasing schemes. From OPES, it inherits a fast-converging bias potential that allows sampling of the relevant part of the phase space. Invernizzi and Parrinello (2022) This facilitates the reweighting, thus guaranteeing a more accurate estimate of the Boltzmann weights and significantly reduces the number of iterations needed for convergence (see Sec. III.2).
On the other hand, from the Kolmogorov bias, we obtain an extensive sampling of the TSE, which, in turn, makes it possible to apply accurately the variational principle (see Sec. III.1).
III.4 Summary of self-consistent iterative procedure
Here, we summarize schematically the whole iterative procedure, in which we alternate training and sampling cycles.
  • Step 1: The NN-based committor $q_{\theta}^{n}(\mathbf{x})$ at iteration $n$ is optimized on the dataset of configurations $\mathbf{x}^{n}$ and weights $w_{i}^{n}$ updated from the previous iterations. To start the procedure, at the first iteration $n=0$, we build such a dataset by running some (short) simulations in the metastable basins and labeling them accordingly. The optimization is based on the variational principle of Eq. 5, minimizing the loss function of Eqs. 8 and 9 (see Sec. III.2).
  • Step 2: We perform biased simulations under the coupled action of the Kolmogorov and OPES bias potentials, $V^{n}_{K}$ and $V^{n}_{\text{OPES}}$. The latter is applied along the committor-related CV $z^{n}(\mathbf{x})$ and drives the sampling across the whole relevant configurational space between the two states, whereas the former guarantees a thorough sampling of the TS region crucial for applying Eq. 5 (see Sec. III.3).
  • Step 3: The dataset is updated with the new sampled configurations, which are reweighted by the effective bias $V_{\text{eff}}=V_{K}+V_{\text{OPES}}$ according to $w^{n}_{i}=\frac{e^{\beta V_{\text{eff}}^{n}(\mathbf{x}_{i})}}{\langle e^{\beta V_{\text{eff}}^{n}(\mathbf{x})}\rangle_{U^{n}_{\text{eff}}}}$. A progressively more accurate estimate of the system’s FES is obtained from such reweighted data.
IV Conclusions
In this work, we have presented a new enhanced sampling strategy for addressing the rare event problem that combines accurate free energy calculations with extensive sampling of transition states. Building on our recent work, Kang, Trizio, and Parrinello (2024) our approach is based on a semi-automatic self-consistent iterative procedure for the determination of the committor function using the variational principle it obeys. A crucial addition to our previous workflow is the definition of a committor-related collective variable that is used not only to drive metadynamics-like calculations but also as an analysis tool for the interpretation of the results to obtain precious physical insights. These improvements lead to much more efficient and accurate learning of the committor function and also to accurate estimates of free energy profiles.
Nonetheless, there is still room for improvement. For example, it would be highly desirable to remove the need for physical descriptors, thus providing an even more general approach, for instance, by exploiting more sophisticated models based on graph neural networks that recently have been making their way also in the enhanced sampling field. Zhang et al. (2024); Dietrich et al. (2024); Zou and Tiwary (2024) On the other hand, from the theoretical point of view, improvements could be obtained by eventually extending the ideas behind the committor function to a more general framework that could help tackle more efficiently complex systems characterized by rugged energetic landscapes and multiple intermediates.
We believe that this approach opens the way to a new era of enhanced sampling, in which reactive processes are studied through the lens of probability distributions, which the committor encodes. Indeed, more and more studies on complex systems are recognizing the importance of embracing the statistical complexity of reactive events in fields as different as catalysis Wang et al. (2015, 2016); Yang, Raucci, and Parrinello (2023); Tian et al. (2023); Perego et al. (2024); Tripathi et al. (2024) and biophysics. Lindorff-Larsen et al. (2011); Dill and MacCallum (2012); Sztain et al. (2021) This further convinces us that a probabilistic description is the natural way to accurately study and analyze in-depth complex reactive processes.
Code and Data Availability
The code for the training of the committor is available through the open-source mlcolvar library Bonati et al. (2023) alongside didactic tutorials. Training and simulation data and inputs are available on GitHub [^1] https://github.com/EnricoTrizio/committor_2.0 alongside the PLUMED Tribello et al. (2014); consortium (2019) interface for the application of the bias.
Authors contributions
All the authors equally contributed to the manuscript by conceptualizing the project, developing the theoretical methodologies, and participating in the writing of the manuscript.
Acknowledgements.
The Authors are grateful to Luigi Bonati, Andrea Rizzi, Jintu Zhang, Umberto Raucci, and Francesco Mambretti for the useful discussions and feedback about this manuscript.
Supporting Information
S1 Numerical stability of Kolmogorov bias potential
To improve the numerical stability and efficacy of the Kolmogorov bias, it is convenient to reformulate its calculation (Eq. 6) in terms of the committor-related $z$ collective variable as it allows lifting the numerical problems that may arise from the direct use of $q$ (see Fig. S1). Here, we show that this is formally equivalent to expressing as a functional of $q$. In the main text, we shown that $q$ is obtained applying a sigmoid-like switching function $\sigma$ to $z$
$q(\mathbf{x})=\sigma(z(\mathbf{x}))\quad\text{with}\quad\sigma(z)=\frac{1}{1+e^{-pz}}$ (S1)
where we set $p=3$. Thus, using the chain rule, we can write
$\frac{\partial{q(\mathbf{x})}}{\partial x}=\frac{\partial{\sigma(z)}}{\partial z}\frac{\partial{z}}{\partial x}$ (S2)
After some elementary algebra, we can equivalently rewrite the bias as
$V_{K}=\frac{1}{\beta}\big{[}\log(|\nabla z|^{2})-4\log({1+e^{-pz}})-2pz\big{]}+c$ (S3)
where $c$ is a constant that thus will not affect free energy calculations.
Such a bias is equivalent to the original one but has the advantage that $\frac{\partial z}{\partial x}$ has a much higher numerical resolution in the metastable basins, where $q\sim 0$ and $q\sim 1$.
Figure S1: Visualization of the numerical issue related to the committor function on a toy double-well potential for which an analytical solution can be obtained.
S2 Müller-Brown potential - Additional information
S2.1 Computational details
Simulations details The Müller-Brown potential energy surface, $U(x,y)$, is defined as a function of the Cartesian coordinates $x$ and $y$
$U(x,y)=-k\sum_{i=1}^{4}d_{i}e^{a_{i}(x-x_{i})+b_{i}(x-x_{i})(y-y_{i})+c_{i}(y-y_{i})}$ (S4)
where the constants take the following values, k = 0.15, [$d_{1}$, $d_{2}$, $d_{3}$, $d_{4}$ ]= [-200, -100, -170, 15], [$a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$] = [-1, -1, -6.5, 0.7], [$b_{1}$, $b_{2}$, $b_{3}$, $b_{4}$] = [0, 0, 11, 0.6], [$c_{1}$, $c_{2}$, $c_{3}$, $c_{4}$] = [-10, -10, -6.5, 0.7], [$x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$] = [1, 0, -0.5, -1] and [$y_{1}$, $y_{2}$, $y_{3}$, $y_{4}$] = [0, 0.5, 1.5, 1].
The simulations of the diffusion of an ideal particle of mass 1 have been performed using Langevin dynamics based on the Bussi-Parrinello algorithm Bussi and Parrinello (2007) as implemented in the ves_md_linearexpansion Valsson and Parrinello (2014) module of PLUMED. Tribello et al. (2014); consortium (2019) The damping constant in the Langevin equation was set to 10/time-unit. The time unit was defined arbitrarily and corresponds to 200 timesteps and natural units ($k_{B}T=1$) were used in all the calculations.
Committor model training details To model the committor function $q_{\theta}(\mathbf{x})$ at each iteration, we used the x and y Cartesian coordinates of the diffusing particle as inputs of a neural network (NN) with architecture [2, 32, 32, 1] nodes/layer. For the optimization, we used the ADAM optimizer with an initial learning rate of $10^{-3}$ modulated by an exponential decay with multiplicative factor $\gamma=0.99999$. The training was performed for 5000 epochs in the first iteration and for $\sim$20000 epochs for the others. The $\alpha$ hyperparameter in the loss function was set to 0.1. The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S1 alongside the lowest value obtained for functional $K_{m}$ (e.q., the variational loss term $L_{v}$), which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
To have a direct comparison with the reference numerical result ${K}_{m}=4.18$, which is taken from Ref. 1, the reported ${K}_{m}$ values are computed on the same ideal dataset described in Ref. 1, which was obtained from a homogeneous grid (that is, 200*200 evenly distributed points) in the relevant part of the Cartesian space (that is, -1.4<x<1.1 and -0.25<y<2.0 ) assigned with their analytical Boltzmann weights.
Table S1: Summary of the iterative procedure for Müller-Brown potential.
      Iteration       Dataset size       $K_{m}$ [au]       OPES BARRIER [$k_{B}T$]       $\lambda$       $t_{s}$ [au]       $t_{o}$ [au]
      0       4000       127.2       -       -       2*400000       200
      1       22000       4.25       20       1       2*5000000       500
      2       40000       4.24       20       1       2*5000000       500
S2.2 Additional figures
Figure S2: Sampling at different iterations on the procedure on the Müller-Brown potential. Points obtained starting the simulation in state A are depicted in blue, and those sampled starting from state B are in red. For the relevant part of the phase space, the 0.5 isolines of a numerical estimate of the committor and the committor estimate at that iteration are provided as a yellow and grey line, respectively.
S3 Alanine Dipeptide - Additional information
S3.1 Computational details
Simulations details The alanine dipeptide (Ace-Ala-Nme) simulations in vacuum have been done using the GROMACS-2021.5 Abraham et al. (2015) MD engine patched with PLUMED Tribello et al. (2014); consortium (2019) and the Amber99-SB Salomon-Ferrer, Case, and Walker (2013) force field with a 2 fs timestep. The Langevin dynamics is sampled with damping coefficient $\gamma_{i}=\frac{m_{i}}{\tau-t}$ with $\tau-t=0.05$ ps and target temperature 300 K.
Committor model training details To model the committor function $q_{\theta}(\mathbf{x})$ at each iteration, we used the 45 distances between all heavy atoms as inputs of a neural network (NN) with architecture [45, 32, 32, 1] nodes/layer. For the optimization, we used the ADAM optimizer with an initial learning rate of $10^{-3}$ modulated by an exponential decay with multiplicative factor $\gamma=0.9999$. The training was performed for 5000 epochs in the first iteration and for $\sim$40000 epochs for the others. The $\alpha$ hyperparameter in the loss function was set to 10. The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S2 alongside the lowest value obtained for the functional $K_{m}$, which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
Table S2: Summary of the iterative procedure for Alanine.
      Iteration       Dataset size       $K_{m}$ [au]       OPES BARRIER [kJ/mol]       $\lambda$       $t_{s}$ [ns]       $t_{o}$ [ps]
      0       4000       5614.68       -       -       4       0.4
      1       22000       6.24       30       0.8       2*10       1
      2       22000       1.91       30       0.8       2*10       1
      3       22000       1.08       30       0.8       2*10       1
S3.2 Additional figures
Figure S3: Sampling at different iterations on the procedure on Alanine. Top row: Scatter plot of the sampled points in the $\varphi\psi$ space. Points obtained starting the simulation in state A (left) are depicted in blue, and those sampled starting from state B (right) are in red.. Middle row: Time series of the $\varphi$ value from a trajectory started from state A. Bottom row: Time series of the $\varphi$ value from a trajectory started from state B.
S4 Double path potential - Additional information
S4.1 Computational details
Simulations details The double path potential energy surface, $U(x,y)$, is defined as a function of the Cartesian coordinates $x$ and $y$
$\begin{split}U(x,y)=10\bigg{(}2+\frac{4x^{4}}{3}-2y^{2}+y^{4}+\frac{10x^{2}(y^{2}-1)}{3}\bigg{)}\\ +7\exp\bigg{(}-\frac{(x+0.7)^{2}+(y-0.8)^{2}}{0.4^{2}}\bigg{)}\\ +\exp\bigg{(}-\frac{(x-1)^{2}+(y+0.3)^{2}}{0.4^{2}}\bigg{)}\\ -6\exp\bigg{(}-\frac{(x+1)^{2}+(y+0.6)^{2})}{0.4^{2}}\bigg{)}\\ -2.35906\end{split}$ (S5)
The simulations of the diffusion of an ideal particle of mass 1 have been performed using Langevin dynamics based on the Bussi-Parrinello algorithm Bussi and Parrinello (2007) as implemented in the ves_md_linearexpansion Valsson and Parrinello (2014) module of PLUMED. Tribello et al. (2014); consortium (2019) The damping constant in the Langevin equation was set to 10/time-unit. The time unit was defined arbitrarily and corresponds to 200 timesteps and natural units ($k_{B}T=1$) were used in all the calculations.
Committor model training details To model the committor function $q_{\theta}(\mathbf{x})$ at each iteration, we used the x and y Cartesian coordinates of the diffusing particle as inputs of a neural network (NN) with architecture [2, 32, 32, 1] nodes/layer. For the optimization, we used the ADAM optimizer with an initial learning rate of $10^{-3}$ modulated by an exponential decay with multiplicative factor $\gamma=0.99999$. The training was performed for 5000 epochs in the first iteration and for $\sim$20000 epochs for the others. The $\alpha$ hyperparameter in the loss function was set to 10. The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S3 alongside the lowest value obtained for functional $K_{m}$ (e.q., the variational loss term $L_{v}$), which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
Table S3: Summary of the iterative procedure for double path potential.
      Iteration       Dataset size       $K_{m}$ [au]       OPES BARRIER [$k_{B}T$]       $\lambda$       $t_{s}$ [au]       $t_{o}$ [au]
      0       4000       35.77       -       -       2*5000000       500
      1       22000       1.38       20       1       2*5000000       500
      2       40000       1.37       20       1       2*5000000       500
S4.2 Additional figures
Figure S4: Sampling at different iterations of the procedure on the double path potential. Points obtained starting the simulation in state A (left) are depicted in blue, and those sampled starting from state B (right) are in red. The 0.5 value of the committor estimate at that iteration is provided as a black line.
S5 Chignolin - Additional information
S5.1 Computational details
Simulations details For the study of folding and unfolding of chignolin (CLN025 peptide sequence Tyr-Tyr-Asp-Pro-Glu-Thr-Gly-Thr-Trp-Tyr) in explicit solvent, we performed our simulations using GROMACS v2021.5 Abraham et al. (2015) patched with PLUMED, Tribello et al. (2014); consortium (2019) the CHARMM22 Piana, Lindorff-Larsen, and Shaw (2011) force field, and the solvent has been modeled by the CHARMM TIP3P MacKerell et al. (1998) force field, sharing the same setup used for long unbiased simulations on this system Lindorff-Larsen et al. (2011) to have a direct comparison with those results. For the same reason, we kept the simulation condition consistent with that work. All simulations were performed with an integration time step of 2 fs and sampling NVT ensemble at 340K. Asp, Glu residues, as well as the N- and C-terminal amino acids are simulated in their charged states. The simulation box contains 1,907 water molecules, together with two sodium ions that neutralize the system. The linear constraint solver algorithm is applied to every bond involving H atoms, and electrostatic interactions are computed via the particle mesh Ewald scheme, with a cutoff of 1 nm for all nonbonded interactions.
Committor model training details To model the committor function $q_{\theta}(\mathbf{x})$ at each iteration, we used the 210 distances suggested by Ref. 19 as inputs of a neural network (NN) with architecture [210, 64, 32, 1] nodes/layer. For the optimization, we used the ADAM optimizer with an initial learning rate of $10^{-3}$ modulated by an exponential decay with multiplicative factor $\gamma=0.99995$. The training was performed for $\sim$30000 epochs. The $\alpha$ hyperparameter in the loss function was set to 1. The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S4 alongside the lowest value obtained for the functional $K_{m}$, which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
Table S4: Summary of the iterative procedure for chignolin.
      Iteration       Dataset size       $K_{m}$ [au]       OPES BARRIER [kJ/mol]       $\lambda$       $t_{s}$ [ns]       $t_{o}$ [ps]
      0       20000       27.36       -       -       2*100       10
      1       220000       4.43       20       0.3       4*250       5
      2       220000       3.64       20       0.3       4*250       5
      3       220000       3.79       20       0.3       4*250       5
      4       220000       2.32       20       0.3       4*250       5
      5       220000       2.17       20       0.3       4*250       5
S5.2 Additional results
Comparison between different descriptor sets In Ref. 1, we trained the committor model using as inputs the 45 distances between $\alpha$-C atoms, whereas here, we used the same set of 210 distances identified in Ref. 19. To compare the two descriptor sets, we optimized the committor model on the same dataset using the two sets, obtaining a $K_{m}$ value of 2.17 for the 210 distances and 18.5 for the 45 distances. This result clearly shows how the greater variational flexibility introduced by side-chain information allows for capturing more features of the folding process.
LASSO Analysis To analyze the relevance of the input feature in the trained model, we applied the LASSO regression analysis Novelli et al. (2022) on several bins along $z$ values. The results of this analysis are summarized in Tab. S5, where the 5 most important descriptors for each bin are reported alongside the corresponding bin range.
Table S5: Summary of LASSO analysis for chignolin.
      z range       Top 5 ranking distances
      $-5.5\sim-4.3$       $d_{102-139},d_{107-139},d_{102-141},d_{109-131},d_{78-141}$
      $-4.3\sim-3.2$       $d_{102-139},d_{102-141},d_{107-139},d_{78-141},d_{109-131}$
      $-3.2\sim-2.0$       $d_{36-102},d_{23-145},d_{36-70},d_{7-145},d_{45-147}$
      $-2.0\sim-0.8$       $d_{22-164},d_{37-69},d_{23-164},d_{37-109},d_{43-120}$
      $-0.8\sim 0.3$       $d_{22-164},d_{23-164},d_{36-105},d_{23-145},d_{37-69}$
      $0.3\sim 1.5$       $d_{23-145},d_{43-120},d_{23-106},d_{36-144},d_{44-152}$
      $1.5\sim 2.7$       $d_{43-120},d_{44-152},d_{23-145},d_{37-109},d_{37-66}$
      $2.7\sim 3.8$       $d_{52-86},d_{63-105},d_{31-100},d_{41-141},d_{56-105}$
      $3.8\sim 5.0$       $d_{43-120},d_{31-115},d_{52-86},d_{49-111},d_{41-144}$
      $5.0\sim 6.1$       $d_{34-128},d_{52-94},d_{52-83},d_{1-152},d_{11-165}$
      z range       Top 5 ranking distances
      $-5.5\sim-4.3$       Gly7-CA:Trp9-CZ2, Thr8-N:Trp9-CZ2, Gly7-CA:Trp9-CH2, Thr8-CA:Trp9-NE1, Glu5-CG:Trp9-CH2
      $-4.3\sim-3.2$       Gly7-CA:Trp9-CZ2, Gly7-CA:Trp9-CH2, Thr8-N:Trp9-CZ2, Glu5-CG:Trp9-CH2, Thr8-CA:Trp9-NE1
      $-3.2\sim-2.0$       Tyr2-CZ:Gly7-CA, Tyr1-O:Tyr10-N, Tyr2-CZ:Pro4-O, Tyr1-CB:Tyr10-N, Asp3-N:Tyr10-CA
      $-2.0\sim-0.8$       Tyr1-C:Tyr10-C, Tyr2-OH:Pro4-C, Tyr1-O:Tyr10-C, Tyr2-OH:Thr8-CA, Tyr2-C:Thr8-O
      $-0.8\sim 0.3$       Tyr1-C:Tyr10-C, Tyr1-O:Tyr10-C, Tyr2-CZ:Gly7-C, Tyr1-O:Tyr10-N, Tyr2-OH:Pro4-C
      $0.3\sim 1.5$       Tyr1-O:Tyr10-N, Tyr2-C:Thr8-O, Tyr1-O:Gly7-O, Tyr2-CZ:Trp9-O, Tyr2-O:Tyr10-CG
      $1.5\sim 2.7$       Tyr2-C:Thr8-O, Tyr2-O:Tyr10-CG, Tyr1-O:Tyr10-N, Tyr2-OH:Thr8-CA, Tyr2-OH:Pro4-CG
      $2.7\sim 3.8$       Asp3-CG:Thr6-N, Pro4-CB:Gly7-C, Tyr2-CG:Gly7-N, Tyr2-CE2:Trp9-CH2, Asp3-O:Gly7-C
      $3.8\sim 5.0$       Tyr2-C:Thr8-O, Tyr2-CG:Thr8-CG2, Asp3-CG:Thr6-N, Asp3-CB:Thr8-CB, Tyr2-CE2:Trp9-O
      $5.0\sim 6.1$       Tyr2-CE1:Trp9-CG, Asp3-CG:Thr6-CG2, Asp3-CG:Glu5-OE2, Tyr1-N:Tyr10-CG, Tyr1-CD1:Tyr10-HE2
Free energy surfaces
Figure S5: Free energy estimates for chignolin projected along $z$ and $q$ obtained averaging on 6 independent simulations: 3 started from the folded state and 3 from the unfolded one. The average value and the relative standard deviation are reported as a solid line and a shaded region, respectively.
S6 Calixarene - Additional information
S6.1 Computational details
Simulations details We performed the simulations with GROMACS Abraham et al. (2015) v2021.5 in combination with the PLUMED Tribello et al. (2014); consortium (2019) plugin. We use the GAFF Wang et al. (2004) force field with RESP Bayly et al. (1993) charges and the TIP3P Jorgensen et al. (1983) water model. We set the timestep to 2 fs and the temperature at 300 K via a velocity rescale thermostat Bussi, Donadio, and Parrinello (2007) with a time constant of 0.1 ps. The simulation box is cubic with a side of 40.27 $\mathrm{\SIUnitSymbolAngstrom}$ and it contains 2100 water molecules in solution together with the host OAMe and the chosen guest molecule. Sodium ions are included to counterbalance excess charges. At every simulation step, the coordinates are aligned so that the vertical axis of the box coincides with the binding axis $h$, and the simulation box is centered on the virtual atom V1.
The funnel restraint In our simulations, we used a funnel restraint Limongelli, Bonomi, and Parrinello (2013) equivalent to the one previously employed by Refs.37; 67; 68 on the same system. Here, we summarize the details of such a restraint, while more details can be found in the original works. The funnel limits the space available to the ligand in the unbound state U by confining it to a cylindrical volume above the binding site. As the ligand approaches the binding site, the funnel restraint becomes wider so that its presence does not affect the binding process itself. Having aligned with PLUMED the system to a reference configuration where the binding axis is found along the vertical axis, we define $h$ as the projection on the binding axis of the center of the carbon atoms of each ligand and $r$ as its radial component. When $h>10$\mathrm{\SIUnitSymbolAngstrom}$$, the funnel surface is a cylinder with radius $R_{cyl}=2$\mathrm{\SIUnitSymbolAngstrom}$$ with its axis along the vertical direction. When $h<10$\mathrm{\SIUnitSymbolAngstrom}$$, the funnel opens into an umbrella-like shape with a 45 degree angle whose surface is defined by $r=12-h$. The force that, for a displacement x, pushes the ligand away from the funnel’s surface is harmonic $-k_{F}x$ with $kF=20$ kJ/mol Å$^{-2}$. A further harmonic restraint is applied on $h$ to prevent the ligand from getting too far from the host, reaching the upper boundary of the simulation box. The corresponding force is $-k_{U}(h-18)$ for $h>18$\mathrm{\SIUnitSymbolAngstrom}$$ and $k_{U}=40$ kJ/mol Å$^{-2}$.
During training, we set boundaries further to state U so that the labeled configurations used to impose the boundary conditions will not influence the committor training in the subsequent iterative simulations. We apply the funnel restraint described above and two additional harmonic restraints $-k_{U}(h-20)$ for $h>20$\mathrm{\SIUnitSymbolAngstrom}$$ and $-k_{U}(h-18)$ for $h<18$\mathrm{\SIUnitSymbolAngstrom}$$, with $k_{U}=20$ kJ/mol.
Because of the funnel presence, the free energy difference between the bound and the true unbound state that we extract from enhanced sampling simulations needs a correction that can be calculated as:
[…]
where $\beta$ = 1/($k_{B}$), $C_{0}$ = 1/1660 Å$^{-3}$ is the standard concentration, $h$ is the coordinate along the funnel’s axis, $W(h)$ is the free energy along the funnel axis and $W_{U}$ its reference value in state U. More precisely, we define $W_{U}$ as the average free energy value in the interval 1.6 $\mathrm{\SIUnitSymbolAngstrom}$ < $h$ < 1.8 $\mathrm{\SIUnitSymbolAngstrom}$. The integral is computed over the state B region that we define as 0.3 $\mathrm{\SIUnitSymbolAngstrom}$ < $h$ < 0.8 $\mathrm{\SIUnitSymbolAngstrom}$.
Committor model training details To model the committor function $q_{\theta}(\mathbf{x})$ at each iteration, we used the same water coordination numbers used in Ref. 37, 8 with respect to 2.5Å-spaced virtual atoms (V) along the binding axis and 4 with respect to 4 atoms on the ligand molecule (M) (2 on the ring (1 and 2), 2 on the terminal atoms (3 and 4)), and a set of distances between the same 4 ligand atoms and the binding pocket, instead of the $h$ projection only, as inputs of a neural network (NN) with architecture [16, 32, 32, 1] nodes/layer. For the optimization, we used the ADAM optimizer with an initial learning rate of $10^{-3}$ modulated by an exponential decay with multiplicative factor $\gamma=0.99995$. The training was performed for $\sim$30000 epochs. The $\alpha$ hyperparameter in the loss function was set to 1. In the loss function, we assigned to the virtual points along the binding axis used for the computation of water coordination numbers an infinite mass, i.e., $10^{6}$. The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S6 alongside the lowest value obtained for the functional $K_{m}$, which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
Table S6: Summary of the iterative procedure for calixarene.
      Iteration       Dataset size       $K_{m}$ [au]       OPES BARRIER [kJ/mol]       $\lambda$       $t_{s}$ [ns]       $t_{o}$ [ps]
      0       2000       27400       -       -       2*10       10
      1       22000       137       50       0.6       2*100       10
      2       22000       333       50       0.6       2*100       10
      3       22000       242       50       0.6       2*100       10
      4       22000       5.16       50       0.6       2*100       10
      4       22000       4.56       50       0.6       2*100       10
S6.2 Additional results
Feature relevance analysis
Figure S6: Descriptors ranking for the committor model of calixarene trained using the descriptors listed in the computational details.
Free energy surfaces
Figure S7: Free energy estimates for calixarene projected along the projection of the ligand molecule on the binding axis $h$, $z$ and $q$.
Figure S8: Two-dimensional free energy surface for calixarene projected in the plane defined by the water coordination number of a virtual point inside the binding cavity (V2) and the projection of the ligand molecule on the binding axis (h).
Kolmogorov Ensemble
Figure S9: Two-dimensional Kolmogorov Ensemble projection for calixarene projected in the plane defined by the water coordination number of a virtual point inside the binding cavity (V2) and the projection of the ligand molecule on the binding axis (h).

### v4

@@
-Everything everywhere all at once, a probability-based enhanced sampling approach to rare events
+Everything everywhere all at once: a probability-based enhanced sampling approach to rare events
 Abstract
 The problem of studying rare events is central to many areas of computer simulations.
-In a recent paper [Kang, P., et al., Nat.
-Comput.
-Sci.
-4, 451–460, 2024], Kang, Trizio, and Parrinello (2024) we have shown that a powerful way of solving this problem passes through the computation of the committor function, and we have demonstrated how the committor can be iteratively computed in a variational way and the transition state ensemble efficiently sampled.
+We recently proposed an approach to solving this problem that passes through the computation of the committor function, showing how it can be iteratively computed in a variational way while efficiently sampling the transition state ensemble.
 Here, we greatly ameliorate this procedure by combining it with a metadynamics-like enhanced sampling approach in which a logarithmic function of the committor is used as a collective variable.
-This integrated procedure leads to an accurate and balanced sampling of the free energy surface in which transition states and metastable basins are studied with the same thoroughness.
-We also show that our approach can be used in cases in which competing reactive paths are possible and intermediate metastable are encountered.
+This procedure leads to an accurate sampling of the free energy surface in which transition states and metastable basins are studied with the same thoroughness.
+We show that our approach can be used in cases with the possibility of competing reactive paths and metastable intermediates.
 In addition, we demonstrate how physical insights can be obtained from the optimized committor model and the sampled data, thus providing a full characterization of the rare event under study.
-We ascribe the success of this approach to the use of a probability-based description of rare events.
 I Introduction
@@
 (2022)
-In the last few decades, our group has contributed to this collective effort and very recently has proposed a new practical method to solve the rare event problem, which is based on computing the committor function $q(\mathbf{x})$ and using it as a sampling tool, which we summarize in Sec.
-III.1Kang, Trizio, and Parrinello (2024) We recall here that the committor is a function of the atomic coordinates $\mathbf{x}$ which, given two metastable states $A$ and $B$, is defined as the probability that a trajectory started at $\mathbf{x}$ ends in $B$ without having first passed by $A$Weinan and Vanden-Eijnden (2010) The committor function is arguably the most precisely defined way of describing rare events and is a quantity that remains well-defined even if the transition from $A$ to $B$ follows different competing pathways.
+In the last few decades, our group has contributed to this collective effort and very recently has proposed a new practical method to address the rare event problem, which is based on computing the committor function $q(\mathbf{x})$ and using it as a sampling tool, which we summarize in Section IV.1Kang, Trizio, and Parrinello (2024) We recall here that the committor is a function of the atomic coordinates $\mathbf{x}$ which, given two metastable states $A$ and $B$, is defined as the probability that a trajectory started at $\mathbf{x}$ ends in $B$ without having first passed by $A$Weinan and Vanden-Eijnden (2010) The committor function is arguably the most precisely defined way of describing rare events and is a quantity that remains well-defined even if the transition from $A$ to $B$ follows different competing pathways.
 As the committor has so many attractive features, much work has been done to determine it.
@@
 $K[q(\mathbf{x})]=\langle|\nabla_{\mathbf{u}}q(\mathbf{x})|^{2}\rangle_{U(\mathbf{x})}$ (1)
-where the average is over the Boltzmann distribution associated with the interatomic potential $U(\mathbf{x})$ at inverse temperature $\beta$, and $\nabla_{\mathbf{u}}$ denotes the gradient with respect to the mass-weighted coordinates.
+where the average $\langle\cdot\rangle_{U(\mathbf{x})}$ is over the Boltzmann distribution associated with the interatomic potential $U(\mathbf{x})$ at inverse temperature $\beta$, and $\nabla_{\mathbf{u}}$ denotes the gradient with respect to the mass-weighted coordinates.
 In Ref.
-1, we computed the variational minimum of $K[q(\mathbf{x})]$ using a self-consistent iterative procedure that to be started, only requires that the initial and final states are known.
-Key to the success of this approach has been the use of a bias potential that is a functional of the committor
+3, we computed the variational minimum of $K[q(\mathbf{x})]$ using a self-consistent iterative procedure that to be started, only requires that the initial and final states are known.
+Key to the success of this approach has been the use of a bias potential $V_{K}$ that is a functional of the committor
 $V_{K}(\mathbf{x})=-\frac{1}{\beta}\log(|\nabla q(\mathbf{x})|^{2})$ (2)
-which is able to focus sampling on the otherwise elusive transition state region (see Fig.
-1).
+which is able to focus sampling on the otherwise elusive transition state region (see Figure 1).
 This approach has proven to be highly successful in finding and extensively sampling the transition state ensemble (TSE).
@@
 This additional calculation can be performed, for instance, using the data collected during the iterative process to generate an efficient collective variable Bonati, Piccini, and Parrinello (2021); Ray, Trizio, and Parrinello (2023) (CV) and then performing a metadynamics-like calculation using this CV.
-Laio and Parrinello (2002); Invernizzi and Parrinello (2020)
+Laio and Parrinello (2002); Invernizzi and Parrinello (2020); Trizio et al.
+(2024)
 Figure 1: Schematic representation of enhanced sampling framework.
@@
 Indeed, in the basins, where $q(\mathbf{x})$ is either $\approx 0$ ($\mathbf{x}\in A$) or $\approx 1$ ($\mathbf{x}\in B$), different configurations are distinguished only by very tiny numerical variations of $q(\mathbf{x})$, thus leading to numerical difficulties.
-Furthermore, another source of numerical issues comes from the very sharp behavior of $q(\mathbf{x})$ in the transition state region, Khoo, Lu, and Ying (2019) further complicating its use as a CV (see Fig.
-1A and SI, Fig.
-S1).
+Furthermore, another source of numerical issues comes from the very sharp behavior of $q(\mathbf{x})$ in the transition state region, Khoo, Lu, and Ying (2019) further complicating its use as a CV (see Figure 1A and Supplementary Figure S1).
 However, a convenient solution can be found if we recall that in Ref.
-1 $q(\mathbf{x})$ was expressed as a neural network whose output $z(\mathbf{x})$ was fed to a step-like activation function $\sigma$ to facilitate imposing the right functional form (see Fig.
-1A)
+3 $q(\mathbf{x})$ was expressed as a neural network whose output $z(\mathbf{x})$ was fed to a step-like activation function $\sigma$ to facilitate imposing the right functional form (see Figure 1A)
 $q(\mathbf{x})=\sigma\big{(}z(\mathbf{x})\big{)}$ (3)
@@
 3 an effective and semi-automatic way to perform free energy calculations.
-Indeed, we do not wait for the end of the iterative process to perform an additional free energy calculation, but we add to $V_{K}$ an On-the-fly Probability Enhanced Sampling Invernizzi and Parrinello (2020) (OPES) bias that uses $z(\mathbf{x})$ as a CV in the iterative procedure itself and it’s aimed at filling the energy landscape to promote transitions (see Fig.
-1B).
-We shall refer to this combined bias as OPES+$V_{K}$, and we show that this, alongside a few technical improvements, accelerates the convergence of the iterative procedure and leads to a balanced sampling, in which configurations belonging to the metastable states and to the transition state ensemble are equally well sampled, and no additional calculations are needed to obtain a converged free energy estimate (see Fig.
-1B).
+Indeed, we do not wait for the end of the iterative process to perform an additional free energy calculation, but, during the whole iterative procedure, we simultaneously apply the $V_{K}$ potential and an On-the-fly Probability Enhanced Sampling Invernizzi and Parrinello (2020); Trizio et al.
+(2024) (OPES) bias that uses $z(\mathbf{x})$ as a CV that it’s aimed at filling the energy landscape to promote transitions (see Figure 1B).
+We shall refer to this combined bias as OPES+$V_{K}$, and we show that this, alongside a few technical improvements, accelerates the convergence of the iterative procedure and leads to a balanced sampling, in which configurations belonging to the metastable states and to the transition state ensemble are equally well sampled, and no additional calculations are needed to obtain a converged free energy estimate upon reweighting of the sampled configurations by the total bias (see Figure 1B).
 For illustrative purposes, we first demonstrate the workings of our approach on some simple but informative systems like some two-dimensional models and the classical example of alanine dipeptide in a vacuum.
@@
 This demonstrates the generality of our approach, which does not require assuming that there is only one reaction pathway.
-A full description of the methods, a summary of the procedure, and other technical details are provided in Sec.
-III and in the SI.
+A full description of the methods, a summary of the procedure, and other technical details are provided in Section IV and in the Supporting Information.
 II Results
@@
 As in our previous work, Kang, Trizio, and Parrinello (2024) we use the $x$ and $y$ coordinates of the particle as descriptors, thus allowing a fair comparison between the two methods.
-Performing OPES+$V_{K}$ simulations allows covering the whole relevant phase space from the metastable states through the TS, as shown in Fig.
-2A.
-Moreover, to achieve this better sampling, it is not necessary to wait until a converged model is available, as good sampling is obtained even from the very first iteration (see SI, Fig.
-S2) in which the model is not much more than a classifier trained to satisfy the boundary conditions $q({\mathbf{x}}_{A})=0$ and $q({\mathbf{x}}_{B})=1$ (see Sec.
-III.2).
-However, this result should come as little surprise, as machine-learning classifier-like CVs have already proved effective in driving enhanced sampling simulations with OPES.
+Performing OPES+$V_{K}$ simulations allows covering the whole relevant phase space from the metastable states through the TS, as shown in Figure 2A.
+Moreover, to achieve this better sampling, it is not necessary to wait until a converged model is available, as good sampling is obtained even from the very first iteration (see Supplementary Figure S2) in which the model is not much more than a classifier trained to satisfy the boundary conditions $q({\mathbf{x}}_{A})=0$ and $q({\mathbf{x}}_{B})=1$ (see Section IV.2).
+However, this result should not surprise, as machine-learning classifier-like CVs have already proved effective in driving enhanced sampling simulations with OPES.
 Bonati, Rizzi, and Parrinello (2020); Trizio and Parrinello (2021); Ray, Trizio, and Parrinello (2023) In contrast, using only $V_{K}$ in Ref.
-1, the sampling was unbalanced toward the TS region and often heavily dependent on the basin from which the simulation was started.
+3, the sampling was unbalanced toward the TS region and often heavily dependent on the basin from which the simulation was started.
 Another advantage in terms of sampling comes from the action of the Kolmogorov bias $V_{K}$, which favors extensive sampling of the TS region.
-This is evident if one compares the distribution along $z$ of the points sampled with a standard OPES simulation and one performed by combining OPES and $V_{K}$ (see Fig.
-2B).
+This is evident if one compares the distribution along $z$ of the points sampled with a standard OPES simulation and one performed by combining OPES and $V_{K}$ (see Figure 2B).
 From the two distributions, it can be clearly appreciated the different emphasis that the two simulations put on the TS region.
@@
 In addition, a second advantage of this approach is related, besides the quantity of the sampling, to its quality.
-Thanks to the static nature of $V_{K}$ and the robust OPES framework, Invernizzi and Parrinello (2022) which is designed to quickly converge to a quasi-static regime, very reliable weights can be assigned to the sampled points, thus allowing for a proper reweighting to the true Boltzmann distribution, which, again, has several implications.
-On the one side, this allows the recovery of the free energy surface (FES), as we show in Fig.
-2C.
+Thanks to the static nature of $V_{K}$ and the robust OPES framework, Invernizzi and Parrinello (2022); Trizio et al.
+(2024) which is designed to quickly converge to a quasi-static regime, very reliable weights can be assigned to the sampled points, thus allowing for a proper reweighting to the true Boltzmann distribution, which, again, has several implications.
+On the one side, this allows the recovery of the free energy surface (FES), as we show in Figure 2C.
 There, we report the free energy profiles along $z$ obtained by reweighting the same two trajectories considered above, compared with an analytical reference.
@@
 On the other side, having such reliable weights also for the points in the metastable regions facilitates the convergence of the iterative procedure, at variance with Ref.
-1.
-Indeed, even for this simple example, this approach allows converging the estimate of the committor to the numerical reference almost perfectly in just a couple of iterations (see inset of Fig.
-2A).
+3.
+Indeed, even for this simple example, this approach allows converging the estimate of the committor to the numerical reference almost perfectly in just two iterations (see inset of Figure 2A).
 II.2 Alanine Dipeptide
 In most publications devoted to rare event sampling, it is traditional to illustrate the capability of a new method on the conformational equilibrium of alanine dipeptide, and here, we shall stick to this tradition.
-Moreover, having studied alanine also in our previous work, Kang, Trizio, and Parrinello (2024) we are in the position to asses the improvement in convergence speed due to our improved protocol.
-There, a good guess for the committor function was obtained from the $6^{\text{th}}$ iteration of the procedure, whereas here, it was already available with half of the iterations (see SI, Table S2).
-This allowed us not only to efficiently sample the whole relevant phase space (see Fig.
-3A) but also to properly identify the TSE (Fig.
-3B), which is known to exhibit a linear relation between $\varphi$ and $\theta$.
+Moreover, having studied alanine also in our previous work, Kang, Trizio, and Parrinello (2024) we are in the position to assess the improvement in convergence speed due to our improved protocol.
+There, a good guess for the committor function was obtained from the $6^{\text{th}}$ iteration of the procedure, whereas here, it was already available with half of the iterations (see Supplementary Table S2).
+This allowed us not only to efficiently sample the whole relevant phase space (see Figure 3A) but also to properly identify the TSE (Figure 3B), which is known to exhibit a linear relation between $\varphi$ and $\theta$.
 Bolhuis, Dellago, and Chandler (2000) It is also worth noting that one can still estimate the Kolmogorov distribution $p_{K}$ by reweighting the OPES+$V_{K}(\mathbf{x})$ sampled configurations only with the OPES part of the bias.
-As discussed in Ref.1, this allows defining in a precise way the TSE.
+As discussed in Ref.3, this allows defining in a precise way the TSE.
 Also in this case, it is instructive to compare the distribution of points sampled using different bias potentials.
-In Fig.
-3C, we compare the outcome of three different OPES simulations, in one, we use the traditional $\varphi$ and $\psi$ as CVs, in the second the $z(\mathbf{x})$ is used as CV, and finally in the third the action of the $z(\mathbf{x})$ OPES bias is boosted by the Kolmogorov bias $V_{K}(\mathbf{x})$.
+In Figure 3C, we compare the outcome of three different OPES simulations, in one, we use the traditional $\varphi$ and $\psi$ as CVs, in the second the $z(\mathbf{x})$ is used as CV, and finally in the third the action of the $z(\mathbf{x})$ OPES bias is boosted by the Kolmogorov bias $V_{K}(\mathbf{x})$.
 As seen for the Müller potential, combining the $V_{K}$ potential with OPES systematically enhances the sampling of the TSE while still guaranteeing a thorough sampling of the metastable states.
-From such data, we can thus recover a correct estimate of the free energy of the system, which is indistinguishable not only from that obtained in the normal OPES run on $z$ but also from the reference simulation using $\varphi$ and $\psi$ (see Fig.
-3D), thus increasing our confidence in the method.
+From such data, we can thus correctly estimate the free energy of the system, which is indistinguishable not only from that obtained in the normal OPES run on $z$ but also from the reference simulation using $\varphi$ and $\psi$ (see Figure 3D), thus increasing our confidence in the method.
 II.3 Double Path Potential
@@
 Another situation that can occur in practice and which is not straightforward to deal with using other methodologies is the one in which the reaction can follow different pathways Vlugt and Smit (2001); Borrero and Dellago (2016); Bolhuis and Csányi (2018); Mandelli, Hirshberg, and Parrinello (2020); Capelli, Carloni, and Parrinello (2019).
-A simplified and yet instructive model is the one depicted in Fig.
-4, where the system can go from state $A$ to state $B$ following two different pathways, with the upper path characterized by a larger free energy barrier.
-With our approach, both paths are sampled starting from the first iteration, thus allowing for a quick convergence of the procedure (see SI, Fig.
-S4).
-It is then instructive to focus on the final results reported in Fig.
-4.
+A simplified and yet instructive model is the one depicted in Figure 4, where the system can go from state $A$ to state $B$ following two different pathways, with the upper path characterized by a larger free energy barrier.
+With our approach, both paths are sampled starting from the first iteration, thus allowing for a quick convergence of the procedure (see Supplementary Figure S4).
+It is then instructive to focus on the final results reported in Figure 4.
 In panel A, we report the converged committor function for the system, showing not only that both paths can be handled at the same time but also that the committor varies more abruptly in the case of the higher barrier, whereas it is more spread for the favored channel.
@@
 7) to identify TSE is to be preferred over the conventional $q\simeq 0.5$ definition.
-If one considers the $p_{K}$ for this system (see Fig.
-4), one can readily see that most of the contribution to the transition flux comes from the lower channel, and only a minor portion goes through the other.
-In contrast, the $q\simeq 0.5$ criterion would have given the same importance to the two paths and also to much higher energy regions that were actually never sampled, not even in biased simulations (see SI, Fig.
-S4).
-The predominance of the lower path is also reflected by the distribution of the sampled points, which are much denser there (see SI, Fig.
-S4).
-Nevertheless, after reweighting, an almost perfect 2D FES for the system, including precise estimates of the barriers, can be recovered effortlessly, as shown in Fig.
-4C.
+If one considers the $p_{K}$ for this system (see Figure 4), one can readily see that most of the contribution to the transition flux comes from the lower channel, and only a minor portion goes through the other.
+In contrast, the $q\simeq 0.5$ criterion would have given the same importance to the two paths and also to much higher energy regions that were actually never sampled, not even in biased simulations (see Supplementary Figure S4).
+The predominance of the lower path is also reflected by the distribution of the sampled points, which are much denser there (see Supplementary Figure S4).
+Eventually, in more complex cases, a more homogeneous sampling could also be achieved by biasing additional orthogonal CVs that could distinguish them, which is also straightforward in the OPES framework (see SectionIV.3).
+Nevertheless, in this case, after reweighting, an almost perfect 2D FES for the system, including precise estimates of the barriers, can be recovered effortlessly, as shown in Figure 4C.
 It is worth noticing that all these results were obtained without any prior information on the system if not knowledge of the initial and final metastable states.
@@
 As discussed in Ref.
-1, better variational flexibility typically leads to more accurate results and to a lower value of the variational bound $K_{m}$.
-Here, in fact, on the same dataset, the minimum goes from the $K_{m}\approx 19$ to the much lower bound of $K_{m}\approx 2.2$ (see also SI, Sec S4).
+3, better variational flexibility typically leads to more accurate results and to a lower value of the variational bound $K_{m}$.
+Here, in fact, on the same dataset, the minimum goes from the $K_{m}\approx 19$ to the much lower bound of $K_{m}\approx 2.2$ (see also Supplementary Section S4).
 […]
 In Ref.
-1, we have shown that the properties of the transition state ensemble could be analyzed in great and illuminating detail with a combination of k-medoid clustering Schubert and Lenssen (2022) and feature relevance analysis.
+3, we have shown that the properties of the transition state ensemble could be analyzed in great and illuminating detail with a combination of k-medoid clustering Schubert and Lenssen (2022) and feature relevance analysis.
 Pizarroso, Portela, and Muñoz (2022) Here, as we get a thorough sampling of the entire reaction path, we extend this analysis to understand the full folding mechanism.
 To do so, we take advantage of the fact that $q(\mathbf{x})$, or equivalently, for simplicity, the smoother $z(\mathbf{x})$, provide a natural ordering in the configurations sampled, that can be used to order them as one goes from $q(\mathbf{x}\in A)=0$ to $q(\mathbf{x}\in B)=1$.
-Thus, we sliced the whole trajectory into 10 bins along the whole $z$ range (i.e., $-6.5<z<6.5$) and into 10 more focused on the TS region (i.e., $-1.5<z<1.5$), and for each, we identified the most representative configurations through the k-medoid analysis.
+Thus, we sliced the whole trajectory into 10 bins along the whole $z$ range (that is, $-6.5<z<6.5$) and into 10 more focused on the TS region (that is, $-1.5<z<1.5$), and for each, we identified the most representative configurations through the k-medoid analysis.
 Schubert and Lenssen (2022) Then, we performed a LASSO-based Novelli et al.
 (2022) feature relevance analysis to identify which descriptors were more relevant for each bin, revealing that the orientation of the side chains does play a role in the formation of the hydrogen bond network that holds the protein together and regulate steric effects.
-In Fig.
-5, we show the superimposition of several configurations belonging to the most representative bins across the (un)folding process.
+In Figure 5, we show the superimposition of several configurations belonging to the most representative bins across the (un)folding process.
 A first general observation is that, as the folding proceeds, the entropy within the bins decreases, and the corresponding configurations are more similar to each other.
-Interestingly, an explanation can be found guided by the LASSO results (see SI, Table S5), which focus the attention on the sidechains of residues Tyr2 and Trp9 in the folded configurations (see Fig.
-5 1 and 2).
+Interestingly, an explanation can be found guided by the LASSO results (see Supplementary Table S5), which focus the attention on the sidechains of residues Tyr2 and Trp9 in the folded configurations (see Figure 5 1 and 2).
 Indeed, moving from frame 1 to 2, the stable folded structure starts to be dismantled due to the misalignment of the side groups of Tyr2 and Trp9, which destroys the hydrophobic $\pi$-$\pi$ interaction between their aromatic rings and owing to the rotation of the side chain of Tyr1, which induces the breaking of the H-bond that locks the end of the main chain.
@@
 […]
-Moreover, our committor-based simulations allowed us to characterize different reaction channels in a single calculation, which we depict in Fig.
-6 and can be identified by visualizing the projection of the Kolmogorov ensemble in the plane defined by $h$ and the water coordination number inside the cavity $V_{2}$ (see SI Fig.
-S9).
-As one could expect, the TS configurations are mostly determined by the distance of the ligand from the host molecule, as proved also by the feature relevance analysis (see SI, Fig.
-S6), which is sensibly shortened with respect to the fully unbound state.
-However, from $p_{K}$, it can also be seen that there are two contributions to the TSE, differing in the behavior of water and, similar to what is discussed in Sec.
-II.3, one dominates over the other.
+Moreover, our committor-based simulations allowed us to characterize different reaction channels in a single calculation, which we depict in Figure 6 and can be identified by visualizing the projection of the Kolmogorov ensemble in the plane defined by $h$ and the water coordination number inside the cavity $V_{2}$ (see Supplementary Figure S9).
+As one could expect, the TS configurations are mostly determined by the distance of the ligand from the host molecule, as proved also by the feature relevance analysis (see Supplementary Figure S6), which is sensibly shortened with respect to the fully unbound state.
+However, from $p_{K}$, it can also be seen that there are two contributions to the TSE, differing in the behavior of water and, similar to what is discussed in Section II.3, one dominates over the other.
 […]
 Figure 7: Convergence of $\boldsymbol{\Delta}\mathbf{G}$ Convergence with simulation time of the estimate for the folding energy of chignolin (A) and the binding energy of the G2 ligand to the OAMe octa-acid guest (B).
-The average estimates from 3 independent simulations for each system are reported as a blue solid line, whereas the uncertainty, computed as the standard deviation over the three replicas, is depicted as a shaded blue region.
+The average estimates from independent simulations for each system (3 for chignolin, 4 for OAMe-G2) are reported as a blue solid line, whereas the uncertainty, computed as the standard deviation over the three replicas, is depicted as a shaded blue region.
 The reference values are provided as grey dashed lines, and the $0.5\,k_{B}T$ interval around the reference is marked by grey dotted lines.
 For chignolin, that is the unbiased estimate from Ref.
-32, for the calixarene, an estimate obtained using the enhanced sampling setup of Ref.
-37.
-III Methods
-In the following sections, we first provide a general background on the committor and the variational principle it obeys, and we summarize the concepts of the Kolmogorov bias and the related ensemble we have introduced in Ref.1 (Sec.
-III.1).
-We then discuss the improved iterative protocol for the self-consistent determination of the committor function (Sec.
-III.2) and its use to converge free energy calculations (Sec.
-III.3), highlighting the advancements from our previous work.
-Kang, Trizio, and Parrinello (2024) In the last section, we schematically summarize the whole iterative procedure to help contextualize all the details (Sec.
-III.4).
-III.1 Background
+33, for the calixarene, an estimate obtained using the enhanced sampling setup of Ref.
+38.
+III Discussion
+In this work, we have presented a new enhanced sampling strategy for addressing the rare event problem that combines accurate free energy calculations with extensive sampling of transition states.
+Building on our recent work, Kang, Trizio, and Parrinello (2024) our approach is based on a semi-automatic self-consistent iterative procedure for the determination of the committor function using the variational principle it obeys.
+A crucial addition to our previous workflow is the definition of a committor-related collective variable that is used not only to drive metadynamics-like calculations but also as an analysis tool for the interpretation of the results to obtain precious physical insights.
+These improvements lead to much more efficient and accurate learning of the committor function and also to accurate estimates of free energy profiles.
+Nonetheless, there is still room for improvement.
+For example, it would be highly desirable to remove the need for physical descriptors, thus providing an even more general approach, for instance, by exploiting more sophisticated models based on graph neural networks that recently have been making their way also in the enhanced sampling field.
+Zhang et al.
+(2024); Dietrich et al.
+(2024); Zou and Tiwary (2024) On the other hand, from the theoretical point of view, improvements could be obtained by eventually extending the ideas behind the committor function to a more general framework that could help tackle more efficiently complex systems characterized by rugged energetic landscapes and multiple intermediates.
+This should also include the possibility of estimating kinetic rates directly from the committor as prescribed in transition path theory.Weinan and Vanden-Eijnden (2010) In addition, it would also be interesting to investigate the applicability and possible advantages of employing our committor-related CV with other enhanced sampling methods in which directly using the committor was problematic.
+Rotskoff, Mitchell, and Vanden-Eijnden (2022); Aristoff et al.
+(2023)
+We believe that this approach opens the way to a new era of enhanced sampling, in which reactive processes are studied through the lens of probability distributions, which the committor encodes.
+Indeed, more and more studies on complex systems are recognizing the importance of embracing the statistical complexity of reactive events in fields as different as catalysis Wang et al.
+(2015, 2016); Yang, Raucci, and Parrinello (2023); Tian et al.
+(2023); Perego et al.
+(2024); Tripathi et al.
+(2024) and biophysics.
+Lindorff-Larsen et al.
+(2011); Dill and MacCallum (2012); Sztain et al.
+(2021) This further convinces us that a probabilistic description is the natural way to accurately study and analyze in-depth complex reactive processes.
+IV Methods
+In the following sections, we first provide a general background on the committor and the variational principle it obeys, and we summarize the concepts of the Kolmogorov bias and the related ensemble we have introduced in Ref.3 (Section IV.1).
+We then discuss the improved iterative protocol for the self-consistent determination of the committor function (Section IV.2) and its use to converge free energy calculations (Section IV.3), highlighting the advancements from our previous work.
+Kang, Trizio, and Parrinello (2024) In the last section, we schematically summarize the whole iterative procedure to help contextualize all the details (Section IV.4).
+IV.1 Background
 The committor function The committor function $q(\mathbf{x})$ was introduced by Kolmogorov for the study of rare events involving transitions between two metastable states $A$ and $B$Kolmogoroff (1931) Such a function is defined as the probability that a trajectory started at configuration $\mathbf{x}$ ends up in $B$ before first passing by $A$ Weinan and Vanden-Eijnden (2010) and is thus believed to be the best possible formal description for a rare event process.
@@
 However, even this approach still presents some serious challenges.
-In a rare event scenario, the committor is indeed almost constant in the metastable basins, i.e., $q\simeq 0$ in A and $q\simeq 1$ in B, and smoothly interpolates between these two values in the transition region.
+In a rare event scenario, the committor is indeed almost constant in the metastable basins, that is, $q\simeq 0$ in A and $q\simeq 1$ in B, and smoothly interpolates between these two values in the transition region.
 As a consequence, the gradients that appear in Eq.
@@
 To bypass this sampling difficulty, in Ref.
-1, we introduced the Kolmogorov bias potential, a new enhanced sampling protocol that, by design, focuses the sampling on the TS, which we introduce in the following section.
+3, we introduced the Kolmogorov bias potential, a new enhanced sampling protocol that, by design, focuses the sampling on the TS, which we introduce in the following section.
 Sampling the Kolmogorov ensemble In a rare event scenario, metastable states are separated by large free energy barriers that hinder the transitions between them.
@@
 Laio and Parrinello (2002); Invernizzi and Parrinello (2020) Indeed, in that case, the bias potential added to the system aims at filling the metastable basins to promote transitions between them by lowering the relative free energy barriers.
-However, in the biased energy landscape, the TS region still remains a local maximum whose sampling is thus unfavored compared to the metastable basins (see Fig.
-1).
+However, in the biased energy landscape, the TS region still remains a local maximum whose sampling is thus unfavored compared to the metastable basins (see Figure 1).
 In addition, the quality of the TSE sampling with such calculations also depends on the quality of the used CV.
@@
 $V_{K}$ depends on the gradients of the committor function, which, as discussed in the previous section, are localized on the TS by construction.
-As a consequence, in the resulting biased energy landscape $U_{K}(\mathbf{x})=U(\mathbf{x})+V_{K}(\mathbf{x})$, the TS region is stabilized over the rest of the phase space, favoring its sampling (see Fig.
-1).
+As a consequence, in the resulting biased energy landscape $U_{K}(\mathbf{x})=U(\mathbf{x})+V_{K}(\mathbf{x})$, the TS region is stabilized over the rest of the phase space, favoring its sampling (see Figure 1).
 Starting from $U_{K}$, we can define the Kolmogorov probability distribution
@@
 Self-consistent approach for the committor In Ref.
-1, we have shown how the variational principle and the sampling of the Kolmogorov ensemble introduced above can be combined in a self-consistent iterative procedure to obtain the committor function with the help of machine learning.
+3, we have shown how the variational principle and the sampling of the Kolmogorov ensemble introduced above can be combined in a self-consistent iterative procedure to obtain the committor function with the help of machine learning.
 In the iterative procedure, we alternate training and sampling cycles until convergence is reached.
@@
 In the following sections, we present an improved version of the procedure that not only allows for a faster convergence through the iterations but also the calculation of the free energy profile of the process.
-III.2 Machine learning the committor function
+IV.2 Machine learning the committor function
 In the method presented in Ref.
-1, the committor function is represented as a neural network (NN) $q_{\theta}(\mathbf{x})=q_{\theta}(\mathbf{d}(\mathbf{x}))$ with learnable parameters $\theta$ that takes as inputs a set of physical descriptors $\mathbf{d}(\mathbf{x})$ (see Fig.
-1).
-Such descriptors are functions of the atomic coordinates and can be, for instance, distances or coordination numbers.
+3, the committor function is represented as a neural network (NN) $q_{\theta}(\mathbf{x})=q_{\theta}(\mathbf{d}(\mathbf{x}))$ with learnable parameters $\theta$ that takes as inputs a set of physical descriptors $\mathbf{d}(\mathbf{x})$ (see Figure 1).
+Such descriptors are functions of the atomic coordinates, for instance, distances or coordination numbers, and as for the case of machine-learning CVs based on the same framework, Bonati et al.
+(2023); Trizio et al.
+(2024) they should be chosen so that they can describe the relevant physics of the systems when considered collectively.
+In addition, in Ref.3, we have shown how the quality of different descriptor sets can be evaluated using the lowest achievable variational bound as a figure of merit.
 The optimization of $q_{\theta}^{n}(\mathbf{x})$, at iteration $n$, is based on the minimization of an objective function $L$ that formalizes the variational principle of Eq.
@@
 In Ref.
-1, the $L_{b}$ term was evaluated on a labeled dataset of $N_{A}$ and $N_{B}$ configurations collected at the first iteration by running short, unbiased MD simulations in the two basins.
+3, the $L_{b}$ term was evaluated on a labeled dataset of $N_{A}$ and $N_{B}$ configurations collected at the first iteration by running short, unbiased MD simulations in the two basins.
 On the other hand, the $L_{v}$ term was computed on the whole dataset available, thus including both the initial unbiased and biased $N^{n}$ configurations from later iterations.
 This was helpful, as the sampling cycles under the action of the $V_{K}$ bias were focused on the TS region but not optimal, as the unbiased configurations were incorrectly weighted in the dataset.
-This issue made it necessary to perform more iterations or correct the corresponding weights (i.e., $w_{i}$ when $i\in A,B$) by an estimate of the free energy difference between the two basins, which is not always readily available.
-Our new approach gets rid of this slowdown, as at each sampling iteration, we can effectively visit not only the TS region but also both metastable states (see Sec.
-III.3).
+This issue made it necessary to perform more iterations or correct the corresponding weights (that is, $w_{i}$ when $i\in A,B$) by an estimate of the free energy difference between the two basins, which is not always readily available.
+Our new approach gets rid of this slowdown, as at each sampling iteration, we can effectively visit not only the TS region but also both metastable states (see Section IV.3).
 In this way, the new configurations cover the whole phase space, and after proper reweighting, the weights $w_{i}$ automatically follow the real Boltzmann distribution.
 As a consequence, we can thus avoid evaluating the $L_{v}$ term on the unbiased labeled configurations, making the whole procedure more precise and faster.
-It should also be noted that, in the first iteration, when only unbiased data are available, the contribution from the $L_{v}$ term was almost null also in the original formulation due to the localization of the gradients $|\nabla q(\mathbf{x})|^{2}$ discussed in Sec.
-III.1.
+It should also be noted that, in the first iteration, when only unbiased data are available, the contribution from the $L_{v}$ term was almost null also in the original formulation due to the localization of the gradients $|\nabla q(\mathbf{x})|^{2}$ discussed in Section IV.1.
 A second consequence of the higher quality of the data thus generated is a more data-efficient protocol.
-Indeed, as soon as the enhanced sampling becomes effective (i.e., multiple transitions can be observed), to optimize the parameters $\theta$ at each iteration, it is sufficient to use configurations and weights generated in the previous iteration or, at most, from the previous two.
-III.3 Extensive sampling along a committor-based CV
+Indeed, as soon as the enhanced sampling becomes effective (that is, multiple transitions can be observed), to optimize the parameters $\theta$ at each iteration, it is sufficient to use configurations and weights generated in the previous iteration or, at most, from the previous two.
+IV.3 Extensive sampling along a committor-based CV
 As already anticipated in the introduction, we bring two conceptual improvements to the original recipe of Ref.
-1.
+3.
 First, we define a physics-informed collective variable (CV) $z(\mathbf{x})$ that is based on the committor and is effective when used with metadynamics-like enhanced sampling approaches.
@@
 A practical way of simply and economically achieving this result comes at zero cost from the NN-based parametrization of the committor proposed in Ref.
-1.
-There, as we have seen in Sec.
-III.2 and depicted in Fig.
-1, the committor is represented as a NN, and a sharp sigmoid activation function $\sigma$ is applied to the output layer of such a network to enforce the correct functional form to the learned function, i.e., $q(\mathbf{x})=\sigma(z(\mathbf{x}))$.
+3.
+There, as we have seen in Section IV.2 and depicted in Figure 1, the committor is represented as a NN, and a sharp sigmoid activation function $\sigma$ is applied to the output layer of such a network to enforce the correct functional form to the learned function, that is, $q(\mathbf{x})=\sigma(z(\mathbf{x}))$.
 It is thus wiser to employ as a CV, not the activated final committor function, but rather the output $z(\mathbf{x})$ of the NN.
@@
 As one can expect, similar numerical issues could also affect the bias potential $V_{K}$ that, in Ref.
-1, was computed as a function of the gradients of the committor function $\nabla q(\mathbf{x})$.
-In fact, this choice is not optimal in practice as, due to the sharp behavior of $q(\mathbf{x})$, such a computation can easily lead to extremely small values below machine precision in the basins where $q(\mathbf{x})$ is flat (see SI, Fig.
-S1), thus making the bias almost ineffective there.
-Conveniently, by taking advantage of the change of variable $q\rightarrow z$ and with the help of some elementary algebra (see SI, Sec.
-S1), one can bypass this issue, writing the bias potential as a function of the smoother and more numerically stable gradients of $z$.
+3, was computed as a function of the gradients of the committor function $\nabla q(\mathbf{x})$.
+In fact, this choice is not optimal in practice as, due to the sharp behavior of $q(\mathbf{x})$, such a computation can easily lead to extremely small values below machine precision in the basins where $q(\mathbf{x})$ is flat (see Supplementary Figure S1), thus making the bias almost ineffective there.
+Conveniently, by taking advantage of the change of variable $q\rightarrow z$ and with the help of some elementary algebra (see Supplementary Section S1), one can bypass this issue, writing the bias potential as a function of the smoother and more numerically stable gradients of $z$.
 Like any CV, $z$ can be used, in principle, with any CV-based method.
-However, here, we opt for the On-the-fly Probability Enhanced Sampling (OPES) method Invernizzi and Parrinello (2020), a recent evolution of metadynamics Laio and Parrinello (2002) that arguably represents the state of the art in this field.
+However, here, we opt for the On-the-fly Probability Enhanced Sampling (OPES) method Invernizzi and Parrinello (2020); Trizio et al.
+(2024), a recent evolution of metadynamics Laio and Parrinello (2002) that arguably represents the state of the art in this field.
 We already pointed out in Ref.
-1 that, in metadynamics-like frameworks, the metastable basins are filled with an external bias potential to reduce the height of free energy barriers and promote transitions, but the TS-region is still found at a local maximum, whose sampling is thus unfavored (see Fig.
-1).
+3 that, in metadynamics-like frameworks, the metastable basins are filled with an external bias potential to reduce the height of free energy barriers and promote transitions, but the TS-region is still found at a local maximum, whose sampling is thus unfavored (see Figure 1).
 Here, we propose to remove this limitation by adding a Kolmogorov bias $V_{K}$ to such a biasing scheme.
@@
 As one can expect, the improved sampling of this new approach is also beneficial to the iterative procedure as it greatly simplifies the dataset construction, exploiting the best from the two biasing schemes.
-From OPES, it inherits a fast-converging bias potential that allows sampling of the relevant part of the phase space.
-Invernizzi and Parrinello (2022) This facilitates the reweighting, thus guaranteeing a more accurate estimate of the Boltzmann weights and significantly reduces the number of iterations needed for convergence (see Sec.
-III.2).
-On the other hand, from the Kolmogorov bias, we obtain an extensive sampling of the TSE, which, in turn, makes it possible to apply accurately the variational principle (see Sec. III.1).
-III.4 Summary of self-consistent iterative procedure
+From the Kolmogorov bias, it inherits an extensive sampling of the TSE, which makes it possible to apply accurately the variational principle (see Section IV.1).
+On the other hand, from OPES, we obtain a fast-converging bias potential that allows sampling of the relevant part of the phase space.
+Invernizzi and Parrinello (2022); Trizio et al.
+(2024) This facilitates the reweighting, thus guaranteeing a more accurate estimate of the Boltzmann weights and substantially reduces the number of iterations needed for convergence (see Section IV.2).
+Moreover, integrating eventual additional CVs in the OPES biasing scheme is simple, which could help to further improve the simulation efficiency in some complex cases, for instance, biasing orthogonal degrees of freedom that distinguish different paths if present.
+IV.4 Summary of self-consistent iterative procedure
 Here, we summarize schematically the whole iterative procedure, in which we alternate training and sampling cycles.
@@
 5, minimizing the loss function of Eqs.
-8 and 9 (see Sec.
-III.2).
+8 and 9 (see Section IV.2).
   • Step 2: We perform biased simulations under the coupled action of the Kolmogorov and OPES bias potentials, $V^{n}_{K}$ and $V^{n}_{\text{OPES}}$.
 The latter is applied along the committor-related CV $z^{n}(\mathbf{x})$ and drives the sampling across the whole relevant configurational space between the two states, whereas the former guarantees a thorough sampling of the TS region crucial for applying Eq.
-5 (see Sec.
-III.3).
+5 (see Section IV.3).
   • Step 3: The dataset is updated with the new sampled configurations, which are reweighted by the effective bias $V_{\text{eff}}=V_{K}+V_{\text{OPES}}$ according to $w^{n}_{i}=\frac{e^{\beta V_{\text{eff}}^{n}(\mathbf{x}_{i})}}{\langle e^{\beta V_{\text{eff}}^{n}(\mathbf{x})}\rangle_{U^{n}_{\text{eff}}}}$.
 A progressively more accurate estimate of the system’s FES is obtained from such reweighted data.
-IV Conclusions
-In this work, we have presented a new enhanced sampling strategy for addressing the rare event problem that combines accurate free energy calculations with extensive sampling of transition states.
-Building on our recent work, Kang, Trizio, and Parrinello (2024) our approach is based on a semi-automatic self-consistent iterative procedure for the determination of the committor function using the variational principle it obeys.
-A crucial addition to our previous workflow is the definition of a committor-related collective variable that is used not only to drive metadynamics-like calculations but also as an analysis tool for the interpretation of the results to obtain precious physical insights.
-These improvements lead to much more efficient and accurate learning of the committor function and also to accurate estimates of free energy profiles.
-Nonetheless, there is still room for improvement.
-For example, it would be highly desirable to remove the need for physical descriptors, thus providing an even more general approach, for instance, by exploiting more sophisticated models based on graph neural networks that recently have been making their way also in the enhanced sampling field.
-Zhang et al.
-(2024); Dietrich et al.
-(2024); Zou and Tiwary (2024) On the other hand, from the theoretical point of view, improvements could be obtained by eventually extending the ideas behind the committor function to a more general framework that could help tackle more efficiently complex systems characterized by rugged energetic landscapes and multiple intermediates.
-We believe that this approach opens the way to a new era of enhanced sampling, in which reactive processes are studied through the lens of probability distributions, which the committor encodes.
-Indeed, more and more studies on complex systems are recognizing the importance of embracing the statistical complexity of reactive events in fields as different as catalysis Wang et al.
-(2015, 2016); Yang, Raucci, and Parrinello (2023); Tian et al.
-(2023); Perego et al.
-(2024); Tripathi et al.
-(2024) and biophysics.
-Lindorff-Larsen et al.
-(2011); Dill and MacCallum (2012); Sztain et al.
-(2021) This further convinces us that a probabilistic description is the natural way to accurately study and analyze in-depth complex reactive processes.
-Code and Data Availability
-The code for the training of the committor is available through the open-source mlcolvar library Bonati et al.
-(2023) alongside didactic tutorials.
-Training and simulation data and inputs are available on GitHub [^1] https://github.com/EnricoTrizio/committor_2.0 alongside the PLUMED Tribello et al.
-(2014); consortium (2019) interface for the application of the bias.
+Data Availability
+Training and simulation data and inputs are available on GitHub tri and on ZenodoTrizio, Kang, and Parrinello (2025)
+Code Availability
+The code for the training of the NN-based committor model alongside didactic tutorials is available through the open-source mlcolvar library Bonati et al.
+(2023), which is the preferred way to access the most updated code.
+To obtain the results reported in the manuscript, version 1.2.2 was used and a frozen version is also available on ZenodoTrizio, Kang, and Parrinello (2025) The PLUMED Tribello et al.
+(2014); consortium (2019) interface for the application of the bias is available on GitHub tri and on ZenodoTrizio, Kang, and Parrinello (2025)
+Acknowledgements.
+The Authors are grateful to Luigi Bonati, Andrea Rizzi, Jintu Zhang, Umberto Raucci, Alice Triveri, and Francesco Mambretti for the useful discussions and feedback about this manuscript.
 Authors contributions
 All the authors equally contributed to the manuscript by conceptualizing the project, developing the theoretical methodologies, and participating in the writing of the manuscript.
-Acknowledgements.
-The Authors are grateful to Luigi Bonati, Andrea Rizzi, Jintu Zhang, Umberto Raucci, and Francesco Mambretti for the useful discussions and feedback about this manuscript.
+Specifically, E.T.
+and P.K.
+developed the code for the training of the committor function and conducted the computational simulations.
+Competing interests statement
+The authors declare no competing interests.
 Supporting Information
@@
 To improve the numerical stability and efficacy of the Kolmogorov bias, it is convenient to reformulate its calculation (Eq.
-6) in terms of the committor-related $z$ collective variable as it allows lifting the numerical problems that may arise from the direct use of $q$ (see Fig.
-S1).
+6) in terms of the committor-related $z$ collective variable as it allows lifting the numerical problems that may arise from the direct use of $q$ (see Figure S1).
 Here, we show that this is formally equivalent to expressing as a functional of $q$.
@@
 Such a bias is equivalent to the original one but has the advantage that $\frac{\partial z}{\partial x}$ has a much higher numerical resolution in the metastable basins, where $q\sim 0$ and $q\sim 1$.
-Figure S1: Visualization of the numerical issue related to the committor function on a toy double-well potential for which an analytical solution can be obtained.
+Figure S1: Numerical issues of the committor function Visualization of the numerical issue related to the committor function on a toy double-well potential along one dimension x for which an analytical solution can be obtained.
 S2 Müller-Brown potential - Additional information
@@
 To have a direct comparison with the reference numerical result ${K}_{m}=4.18$, which is taken from Ref.
-1, the reported ${K}_{m}$ values are computed on the same ideal dataset described in Ref.
-1, which was obtained from a homogeneous grid (that is, 200*200 evenly distributed points) in the relevant part of the Cartesian space (that is, -1.4<x<1.1 and -0.25<y<2.0 ) assigned with their analytical Boltzmann weights.
-Table S1: Summary of the iterative procedure for Müller-Brown potential.
+3, the reported ${K}_{m}$ values are computed on the same ideal dataset described in Ref.
+3, which was obtained from a homogeneous grid (that is, 200*200 evenly distributed points) in the relevant part of the Cartesian space (that is, -1.4<x<1.1 and -0.25<y<2.0 ) assigned with their analytical Boltzmann weights.
+Table S1: Summary Müller-Brown Summary of the iterative procedure for Müller-Brown potential.
       Iteration       Dataset size       $K_{m}$ [au]       OPES BARRIER [$k_{B}T$]       $\lambda$       $t_{s}$ [au]       $t_{o}$ [au]
@@
 The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S2 alongside the lowest value obtained for the functional $K_{m}$, which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
-Table S2: Summary of the iterative procedure for Alanine.
+Table S2: Summary Alanine Summary of the iterative procedure for Alanine.
       Iteration       Dataset size       $K_{m}$ [au]       OPES BARRIER [kJ/mol]       $\lambda$       $t_{s}$ [ns]       $t_{o}$ [ps]
@@
 S3.2 Additional figures
-Figure S3: Sampling at different iterations on the procedure on Alanine.
+Figure S3: Iterations sampling Alanine Sampling at different iterations on the procedure on Alanine.
 Top row: Scatter plot of the sampled points in the $\varphi\psi$ space.
@@
 The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S3 alongside the lowest value obtained for functional $K_{m}$ (e.q., the variational loss term $L_{v}$), which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
-Table S3: Summary of the iterative procedure for double path potential.
+Table S3: Summary double-path potential Summary of the iterative procedure for double path potential.
       Iteration       Dataset size       $K_{m}$ [au]       OPES BARRIER [$k_{B}T$]       $\lambda$       $t_{s}$ [au]       $t_{o}$ [au]
@@
 S4.2 Additional figures
-Figure S4: Sampling at different iterations of the procedure on the double path potential.
+Figure S4: Iterations sampling double-path Sampling at different iterations of the procedure on the double path potential.
 Points obtained starting the simulation in state A (left) are depicted in blue, and those sampled starting from state B (right) are in red.
@@
 The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S4 alongside the lowest value obtained for the functional $K_{m}$, which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
-Table S4: Summary of the iterative procedure for chignolin.
+Table S4: Summary Chignolin Summary of the iterative procedure for chignolin.
       Iteration       Dataset size       $K_{m}$ [au]       OPES BARRIER [kJ/mol]       $\lambda$       $t_{s}$ [ns]       $t_{o}$ [ps]
@@
 Comparison between different descriptor sets In Ref.
-1, we trained the committor model using as inputs the 45 distances between $\alpha$-C atoms, whereas here, we used the same set of 210 distances identified in Ref.
+3, we trained the committor model using as inputs the 45 distances between $\alpha$-C atoms, whereas here, we used the same set of 210 distances identified in Ref.
 19.
@@
 S5, where the 5 most important descriptors for each bin are reported alongside the corresponding bin range.
-Table S5: Summary of LASSO analysis for chignolin.
+Table S5: LASSO summary Chignolin Summary of LASSO analysis for chignolin.
+The atoms involved in the distances used as descriptors are indicated according to the indexes used in PLUMED (upper table) and their names from the PDB (lower table).
       z range       Top 5 ranking distances
@@
 Free energy surfaces
-Figure S5: Free energy estimates for chignolin projected along $z$ and $q$ obtained averaging on 6 independent simulations: 3 started from the folded state and 3 from the unfolded one.
+Figure S5: Free energy surface Chignolin Free energy estimates for Chignolin projected along $z$ and $q$ obtained averaging on 6 independent simulations: 3 started from the folded state and 3 from the unfolded one.
 The average value and the relative standard deviation are reported as a solid line and a shaded region, respectively.
@@
 At every simulation step, the coordinates are aligned so that the vertical axis of the box coincides with the binding axis $h$, and the simulation box is centered on the virtual atom V1.
-The funnel restraint In our simulations, we used a funnel restraint Limongelli, Bonomi, and Parrinello (2013) equivalent to the one previously employed by Refs.37; 67; 68 on the same system.
+The funnel restraint In our simulations, we used a funnel restraint Limongelli, Bonomi, and Parrinello (2013) equivalent to the one previously employed by Refs.38; 70; 71 on the same system.
 Here, we summarize the details of such a restraint, while more details can be found in the original works.
@@
 Committor model training details To model the committor function $q_{\theta}(\mathbf{x})$ at each iteration, we used the same water coordination numbers used in Ref.
-37, 8 with respect to 2.5Å-spaced virtual atoms (V) along the binding axis and 4 with respect to 4 atoms on the ligand molecule (M) (2 on the ring (1 and 2), 2 on the terminal atoms (3 and 4)), and a set of distances between the same 4 ligand atoms and the binding pocket, instead of the $h$ projection only, as inputs of a neural network (NN) with architecture [16, 32, 32, 1] nodes/layer.
+38, 8 with respect to 2.5Å-spaced virtual atoms (V) along the binding axis and 4 with respect to 4 atoms on the ligand molecule (M) (2 on the ring (1 and 2), 2 on the terminal atoms (3 and 4)), and a set of distances between the same 4 ligand atoms and the binding pocket, instead of the $h$ projection only, as inputs of a neural network (NN) with architecture [16, 32, 32, 1] nodes/layer.
 For the optimization, we used the ADAM optimizer with an initial learning rate of $10^{-3}$ modulated by an exponential decay with multiplicative factor $\gamma=0.99995$.
@@
 The $\alpha$ hyperparameter in the loss function was set to 1.
-In the loss function, we assigned to the virtual points along the binding axis used for the computation of water coordination numbers an infinite mass, i.e., $10^{6}$.
+In the loss function, we assigned to the virtual points along the binding axis used for the computation of water coordination numbers an infinite mass, that is, $10^{6}$.
 The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations are summarized in Table S6 alongside the lowest value obtained for the functional $K_{m}$, which provides a quality and convergence measure, the simulation time $t_{s}$ and the output sampling time $t_{o}$.
-Table S6: Summary of the iterative procedure for calixarene.
+Table S6: Summary calixarene Summary of the iterative procedure for calixarene.
       Iteration       Dataset size       $K_{m}$ [au]       OPES BARRIER [kJ/mol]       $\lambda$       $t_{s}$ [ns]       $t_{o}$ [ps]
@@
 Feature relevance analysis
-Figure S6: Descriptors ranking for the committor model of calixarene trained using the descriptors listed in the computational details.
+Figure S6: Feature relevance calixarene Descriptors ranking for the committor model of calixarene trained using the descriptors listed in the computational details.
 Free energy surfaces
-Figure S7: Free energy estimates for calixarene projected along the projection of the ligand molecule on the binding axis $h$, $z$ and $q$.
-Figure S8: Two-dimensional free energy surface for calixarene projected in the plane defined by the water coordination number of a virtual point inside the binding cavity (V2) and the projection of the ligand molecule on the binding axis (h).
+Figure S7: 1D free energy surface calixarene Free energy estimates for calixarene projected along the projection of the ligand molecule on the binding axis $h$, $z$ and $q$.
+Figure S8: 2D free energy surface calixarene Two-dimensional free energy surface for calixarene projected in the plane defined by the water coordination number of a virtual point inside the binding cavity (V2) and the projection of the ligand molecule on the binding axis (h).
 Kolmogorov Ensemble
-Figure S9: Two-dimensional Kolmogorov Ensemble projection for calixarene projected in the plane defined by the water coordination number of a virtual point inside the binding cavity (V2) and the projection of the ligand molecule on the binding axis (h).
+Figure S9: Kolomogorov distribution calixarene Two-dimensional Kolmogorov probability $p_{K}$ projection for calixarene projected in the plane defined by the water coordination number of a virtual point inside the binding cavity (V2) and the projection of the ligand molecule on the binding axis (h).